### v1

# Inhomogeneous nonlinear Schrödinger equations with competing singular nonlinearities

Abstract
We study nonlinear elliptic equations arising as stationary states of inhomogeneous nonlinear Schrödinger equations with competing singular nonlinearities. Working in a weighted Sobolev space that combines the homogeneous Sobolev space with a weighted Lebesgue term, we establish continuous and compact embeddings of Caffarelli–Kohn–Nirenberg type. These embeddings, together with a model that displays a natural scaling, allow us to apply the abstract critical point framework of Mercuri and Perera (2025), yielding a sequence of nonlinear eigenvalues for the associated problem. This scaling property leads to a classification of weighted power-type nonlinearities into subscaled, scaled, and superscaled regimes. Within this variational setting, we obtain broad existence and multiplicity results for equations driven by sums of weighted power nonlinearities, covering superscaled, scaled, and subscaled interactions, both in the subcritical and critical cases. We also provide a nonexistence result as a consequence of a Pohozaev-type identity. Finally, in the radial setting we employ improved radial CKN inequalities to enlarge the admissible embedding ranges. This yields strengthened radial versions of all our main results, including two-dimensional configurations with more singular weights, where no compact embeddings are available in the nonradial case.
Keywords: inhomogeneous nonlinear Schrödinger equations; singular weighted elliptic problems; nonlinear eigenvalues; variational methods.
MSC 2020: Primary: 35J60, 35J20; Secondary: 35Q55, 35B33, 35P30, 58E05.
1 Introduction
Elliptic equations with singular or varying coefficients appear naturally in the study of several physical models such as plasma physics, nonlinear optics, and Bose–Einstein condensation. In these contexts, the local strength of the nonlinearity depends on the spatial position, leading to so-called inhomogeneous nonlinear Schrödinger equations (INLS). The time-dependent INLS
$i\partial_{t}u+\Delta u=\lambda|x|^{-a}|u|^{p-2}u+\mu|x|^{-b}|u|^{q-2}u,\qquad(t,x)\in\mathbb{R}\times\mathbb{R}^{N},$
exhibits rich dynamics involving the competition between two nonlinearities with distinct homogeneities. Depending on the signs of the parameters, one nonlinearity may be focusing and the other defocusing, leading to the coexistence of scattering and blow-up phenomena, as shown for instance in [5, 8, 13, 14, 17, 18, 19, 27]. The search for stationary wave solutions of such evolution problems gives rise to a large family of nonlinear elliptic equations of Schrödinger type with singular or decaying weights.
The study of weighted nonlinear elliptic equations has a long tradition, deeply connected with the development of weighted Sobolev embeddings and Caffarelli–Kohn–Nirenberg (CKN) inequalities [7, 22]. The presence of weights such as $|x|^{-a}$ or $|x|^{-b}$ typically breaks translation and scaling invariance, which requires the analysis to be carried out in suitable weighted Sobolev spaces. Early works addressing these difficulties include Wang and Willem [34], who considered singular minimization problems for the Laplacian with critical weighted nonlinearities that restore scaling invariance and proved existence of solutions via minimization of the best embedding constant. Ambrosetti, Felli and Malchiodi [1] studied Schrödinger equations with potentials vanishing at infinity and established embeddings in weighted Sobolev spaces, obtaining ground state solutions for nonlinearities in the range $\sigma<p<2^{*}$, where $\sigma$ depends on the decay of the potential. Related approaches closer to the problems treated in the present paper were developed by Sintzoff [29, 30] and by Su, Wang and Willem [33], who considered unbounded and decaying radial potentials, establishing compact weighted embeddings and the existence of solutions. These results were later partially extended to the $p$-Laplacian setting by Su and Tian [32] and by Assunção, Carrião and Miyagaki [2], who studied $p$-Laplacian equations with singular weights, employing Caffarelli–Kohn–Nirenberg inequalities and a symmetric Mountain Pass framework to obtain infinitely many solutions. Further developments on eigenvalue-type problems with singular potentials were obtained by Chabrowski and Costa [9], who studied multiplicity and bifurcation near the first eigenvalue. Other contributions along these lines, considering bounded and unbounded domains, include [3, 4, 10, 15, 16, 20, 21, 24, 28].
In parallel, considerable progress has been made in the dispersive framework associated with the time-dependent INLS equation. For the three-dimensional cubic focusing case, Farah and Guzmán [13] obtained scattering of radial solutions below the ground-state threshold, later extending their results to higher dimensions in [14]. Campos [8] and Hajaiej, Luo and Wang [19] refined these techniques for more general exponents and in the presence of inverse-square potentials, while Bellazzini, Dinh and Forcella [5] employed the interaction Morawetz method to treat the nonradial problem with combined nonlinearities. The recent papers by Gou, Majdoub and Saanouni [17, 18] have also studied INLS equations with competing inhomogeneous terms, both in the radial and nonradial settings. These works highlight the role of the associated elliptic stationary equations in determining the threshold dynamics of the corresponding evolution problem.
Motivated by these results, we turn our attention to the elliptic stationary problem,
$\Delta u+\frac{|u|^{q-2}u}{|x|^{b}}=\lambda\,\frac{|u|^{p-2}u}{|x|^{a}}\qquad\text{in }\mathbb{R}^{N},$ (1.1)
where $0<a,b<2$, $2<p<q$, and $N\geq 2$. This problem can be viewed as the search for standing wave profiles of the time-dependent INLS with two competing nonlinearities.
Throughout this work, we define, for any $0\leq\eta<N$,
$2^{*}_{\eta}=\left\{\begin{aligned} &+\infty&\text{ if }&\quad N=2\\ &\frac{2(N-\eta)}{N-2}&\text{ if }&\quad N\geq 3,\end{aligned}\right$.
and we assume $2<q<2^{*}_{b}$. This is equivalent to requiring that N<2(qb)/(q2)N<2(q-b)/(q-2). Let
$\delta=\frac{2-b}{q-2}.$
The associated scaling given by
$u_{t}(x)=t^{\delta}u(tx),\,\,t\in[0,\infty),\,\,x\in\mathbb{R}^{N}$ (1.2)
leaves the equation invariant when the parameters satisfy
$a=a(b,q,p):=2-\delta(p-2)=b+\delta(q-p)\in(b,2).$ (1.3)
If $u$ is a solution of equation (1.1), then the entire 1-parameter family of functions $\{u_{t}:t\geq 0\}$ is a solution. Therefore, (1.1) can be regarded as a nonlinear eigenvalue problem. In this work we exploit the theory developed in [23] to study (1.1) and related problems of the form
$\Delta u+\frac{|u|^{q-2}u}{|x|^{b}}=f(x,u)\quad\text{in }\mathbb{R}^{N}.$ (1.4)
To emphasize the interaction between $f$ and the nonlinear eigenvalues for (1.1), we usually set
$f(x,u)=\lambda\frac{|u|^{p-2}u}{|x|^{a}}+g(x,u)$
where $g$ is a Carathéodory function on $\mathbb{R}^{N}\times\mathbb{R}$ satisfying a suitable growth condition. Throughout most of this paper, we work under the following basic assumptions on the parameters:
[…]
By using sharp weighted embeddings of Caffarelli–Kohn–Nirenberg type and their radial refinements [7, 22], we obtain continuous and compact embeddings of the relevant weighted Sobolev spaces into suitable weighted Lebesgue spaces. These embeddings allow for a variational construction of nonlinear eigenpairs $(\lambda_{k},u_{k})$ for problem (1.1) via a min–max scheme based on the cohomological index of Fadell and Rabinowitz [12]. Our approach treats one of the nonlinear terms as part of the operator, extending the classical frameworks and permitting a significantly broader range of parameters and growth conditions.
Within this setting, we carry out a comprehensive analysis of nonlinearities given by sums of weighted power-type terms, under assumptions dictated by their asymptotic behavior relative to the scaled nonlinearity and the sequence of eigenvalues. We establish existence and multiplicity results for (1.4) across all the corresponding regimes. The precise statements of our main theorems are presented in Subsection 1.2.
1.1 Variational framework
We define the natural energy space to look for solutions to problems (1.1) and (1.4). For $N\geq 2$, $r\geq 1$ and $\eta\in(0,N)$ we denote by $L^{r}_{\eta}(\mathbb{R}^{N})$ the weighted Lebesgue space $L^{r}(\mathbb{R}^{N};|x|^{-\eta}dx)$. Recall that $L^{r}_{\eta}(\mathbb{R}^{N})$ is a Banach space with the norm
$\|u\|_{L^{r}_{\eta}}=\left(\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx\right)^{\frac{1}{r}}$
and it is uniformly convex if $r>1$. Considering $b\in(0,N)$ and $q\geq 1$, we define
$E^{q}_{b}(\mathbb{R}^{N})=\left\{u\in L^{q}_{b}(\mathbb{R}^{N}):\nabla u\in(L^{2}(\mathbb{R}^{N}))^{N}\right\},$
where $\nabla u=(u_{x_{1}},\cdots,u_{x_{N}})$ represents the weak derivatives of $u$, endowed with the norm
$\|u\|$ $=\left[\int_{\mathbb{R}^{N}}|\nabla u|^{2}\,dx+\left(\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx\right)^{\frac{2}{q}}\right]^{\frac{1}{2}}.$
We readily see that $E_{b}^{q}(\mathbb{R}^{N})$ (briefly denoted by $E_{b}^{q}$) is a Banach space. For $N\geq 3$, the space $C_{c}^{\infty}(\mathbb{R}^{N})$} is dense in $E_{b}^{q}(\mathbb{R}^{N})$, while for $N=2$ this density holds provided that $q\geq b$. Since we could not locate a direct reference for this rather standard fact, we include a complete proof in the Appendix (see Proposition 6.1). Moreover, for $q>1$, recalling that the product space $L^{q}_{b}(\mathbb{R}^{N})\times(L^{2}(\mathbb{R}^{N}))^{N}$ is uniformly convex with
$\|(u,v_{1},\cdots,v_{N})\|_{1}=\left(\|u\|_{L^{q}_{b}}^{2}+\sum_{i=1}^{N}\|v_{i}\|_{L^{2}}^{2}\right)^{\frac{1}{2}},$
since the operator $T:E^{q}_{b}\to L^{q}_{b}(\mathbb{R}^{N})\times(L^{2}(\mathbb{R}^{N}))^{N}$ given by $T(u)=(u,\nabla u)$ is an isometry, the Banach space $E^{q}_{b}$ is also uniformly convex. In particular, it is reflexive when $q>1$.
As usual, a key step in formulating the energy functional associated with our problem is to determine into which Lebesgue spaces the underlying Sobolev-type space embeds. For $N\geq 2$, $0<b<2<q<2^{*}_{b}$ and $0\leq\eta<N$ let us define
$2^{*}_{b,q,\eta}:=\begin{cases}\vskip 5.69046pt\frac{q(N-\eta)}{N-b}&if\quad N\geq 2\quad\text{and}\quad b\leq\eta,\\ \frac{\eta\,q\,(N-2)+2N(b-\eta)}{b\,(N-2)}&if\quad N\geq 3\quad\text{and}\quad 0\leq\eta<b.\end{cases}$ (1.5)
Note that $2^{*}_{b,q,\eta}\leq\min\{q,2^{*}_{\eta}\}<2^{*}$ when $b\leq\eta$, while for $0\leq\eta<b$ and $N\geq 3$ one has $q<2^{*}_{b,q,\eta}<2^{*}_{\eta}\leq 2^{*}$.
Under suitable conditions on $\eta$, we shall prove that $E_{b}^{q}$ embeds continuously (and, in some regimes, compactly) into $L_{\eta}^{r}(\mathbb{R}^{N})$ for $r$ lying between $2^{*}_{b,q,\eta}$ and $2^{*}_{\eta}$ with the endpoints included depending on some additional conditions. The precise hypotheses and the corresponding embedding results will be established in Section 2.
Given a Carathéodory function $f:\mathbb{R}^{N}\times\mathbb{R}\to\mathbb{R}$, we say that $u\in E^{q}_{b}$ is a weak solution for the problem
$\Delta u+\frac{|u|^{q-2}u}{|x|^{b}}=f(x,u)\qquad\text{in }\mathbb{R}^{N},$
if $f(\cdot,u)\in L^{1}_{loc}(\mathbb{R}^{N})$ and
$\int_{\mathbb{R}^{N}}\nabla u\nabla vdx+\int_{\mathbb{R}^{N}}\frac{|u|^{q-2}uv}{|x|^{b}}dx=\int_{\mathbb{R}^{N}}f(x,u)vdx,\quad\text{for all}\ v\in C^{\infty}_{c}(\mathbb{R}^{N}).$
Thus, the natural energy functionals associated with problems (1.1) and (1.4) are $\Phi_{\lambda},\Phi:E^{q}_{b}\to\mathbb{R}$ defined by
$\begin{array}[]{c}\displaystyle\Phi_{\lambda}(u)=\frac{1}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\frac{1}{q}\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx-\frac{\lambda}{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx,\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \displaystyle\Phi(u)=\frac{1}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\frac{1}{q}\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx-\int_{\mathbb{R}^{N}}F(x,u)dx,\end{array}$ (1.6)
where $F(x,s)=\int_{0}^{s}f(x,t)\,dt$. The growth assumptions on $F$ will be stated in terms of the embeddings introduced above and will ensure that $\Phi\in C^{1}$. Under the corresponding conditions on the pair $(a,p)$, the same embeddings also guarantee that $\Phi_{\lambda}\in C^{1}$. Therefore, a function $u\in E_{b}^{q}$ is a solution of (1.1) (resp. (1.4)) if and only if it is a critical point of $\Phi_{\lambda}$ (resp. $\Phi$).
1.2 Main theorems
Before stating the main results, we establish some notations and definitions that will be used throughout this paper.
Given the definition of $u_{t}$ presented in (1.2), for $\eta\in[0,N)$ and $r\geq 1$ we have
$\int_{\mathbb{R}^{N}}\frac{|u_{t}|^{r}}{|x|^{\eta}}dx=t^{r\delta+\eta-N}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx,\quad\forall t\geq 0,\,u\in L^{r}_{\eta}(\mathbb{R}^{N}).$
Notice that $r\delta+\eta-N>0$ for any $r\geq 2^{*}_{b,q,\eta}$ (see (1.5)). Since this change of variables is a common argument in this work, from now on we denote
$\ell(\eta,r):=r\delta+\eta-N\qquad\text{and}\qquad\ell:=\ell(b,q)=\ell(a,p).$ (1.7)
Problem (1.1) yields a natural classification of possibly different growth regimes for the nonlinearities that we will consider in problem (1.4).
Definition 1.
Assume ($\mathcal{H}$). We say that the pair $(\eta,r)$ is admissible if
$0\leq\eta<\frac{N+2}{2}\ \ \big(b\leq\eta<2,\text{ if }N=2\big)\quad\text{and}\quad r\in(\,2^{*}_{b,q,\eta},\;2^{*}_{\eta}\,),\quad\text{with}\quad r>1.$
In addition, when $\eta\leq b$ we allow $r=2^{*}_{b,q,\eta}$, and if $\eta\leq 2$ and $N\geq 3$ we include the second endpoint $r=2^{*}_{\eta}$. We classify an admissible pair $(\eta,r)$ in three types:
  1. (i)
    Subscaled, if $\ell(\eta,r)<\ell$;
  2. (ii)
    Scaled, if $\ell(\eta,r)=\ell$;
  3. (iii)
    Superscaled, if $\ell(\eta,r)>\ell$.
The concept of admissible pairs, together with the conditional inclusion of the endpoints of the above intervals, is motivated by the embeddings established in Lemma 2.2 and it also relates with the nonexistence result given in Proposition 2.5. The restriction $r>1$ is necessary for the $C^{1}$ regularity of the functionals associated to the problems.
Notice that if $(\eta,r)$ is subscaled (scaled, superscaled) then
$r<q+\frac{(b-\eta)}{\delta}\quad\left(r=q+\frac{(b-\eta)}{\delta},\quad r>q+\frac{(b-\eta)}{\delta}\right).$
Thus, the existence of a subscaled (scaled) pair $(\eta,r)$ requires $\eta>b$ ($\eta\geq b$), since for $N\geq 3$ one has
$q+\frac{b-\eta}{\delta}<\frac{q(N-\eta)}{N-b}<2^{*}_{b,q,\eta},\quad\text{for all}\quad\eta\in[0,b).$
In other words, when $0\leq\eta<b$, all admissible pairs are necessarily superscaled. On the other hand, the condition $\eta<2$ is necessary for the existence of a superscaled pair, since
$q+\frac{b-\eta}{\delta}\;\geq\;2^{*}_{\eta}\qquad\text{whenever }\eta\geq 2.$
Moreover, although $0\leq\eta<N$ would be a more natural assumption for the singularities, the initial restriction $\eta<(N+2)/2$ in this definition is to ensure that there will always exist a number $r>1$ in the range considered.
We begin by proving existence of nontrivial solutions for subcritical superscaled problems. Our first theorem reads
Theorem 1.1.
Assume ($\mathcal{H}$). Suppose $f(x,t)=\lambda|x|^{-a}|t|^{p-2}t+|x|^{-\eta}h(t)$, with $\lambda\in\mathbb{R}$, $0<\eta<N-q(N-2)/2$ ($\eta\in(b,2)$ if $N=2$) and $h=H^{\prime}$ for $H:\mathbb{R}\to\mathbb{R}$ a $C^{1}$ function such that
$|h(t)|\leq c_{1}|t|^{r_{1}-1}+c_{2}|t|^{r_{2}-1}\quad\text{and}\quad c|t|^{r}\leq rH(t)\leq h(t)t,\quad\forall t\in\mathbb{R},$ (1.8)
for some constants $c,c_{1},c_{2}>0$, and superscaled pairs $(\eta,r_{1})$, $(\eta,r_{2})$ and $(\eta,r)$, with $2^{*}_{b,q,\eta}<r_{1},r_{2}<2^{*}_{\eta}$ and $q<r<2^{*}_{\eta}$. Then, problem (1.4) has a nontrivial solution. Moreover, if $\lambda$ is not an eigenvalue of (1.1), this solution has positive energy.
Theorem 1.2.
Besides the conditions of Theorem 1.1 we assume that $h$ is an odd function. Then, problem (1.4) has a sequence $(u_{k})$ of nontrivial solutions satisfying $\Phi(u_{k})\nearrow\infty$.
In the next results we consider problem (1.4) with critical nonlinearities as
$f(x,u)=\lambda\frac{|u|^{p-2}u}{|x|^{a}}+\frac{|u|^{2^{*}_{\eta_{1}}-2}u}{|x|^{\eta_{1}}}+\mu\frac{|u|^{r_{2}-2}u}{|x|^{\eta_{2}}}$ (1.9)
for $\lambda\in\mathbb{R}$, $\mu\geq 0$ and $N\geq 3$. Suppose that $\eta=\eta_{1}$ satisfies
$0\leq\eta<N-q\frac{N-2}{2}.$ (1.10)
For $(\eta_{2},r_{2})$, we assume that it is a superscaled pair satisfying one of the following conditions:
$\eta_{2}\in[0,\eta_{1})\quad\text{and}\quad 2^{*}_{\eta_{1}}\leq r_{2}<2^{*}_{\eta_{2}}.$ (1.11)
$\eta_{2}\in[\eta_{1},2)\,\,(\eta_{2}=\eta_{1}\,\,\text{only if}\,\,\eta_{1}>b)\quad\text{and}\quad r_{2}<2^{*}_{\eta_{2}},\quad\text{with}\,\,r_{2}>q+\frac{(b-\eta_{2})(r_{1}-q)}{b-\eta_{1}}\quad\text{if}\quad\eta_{2}<b.$ (1.12)
[…]
These results depend on the position of $\lambda$ relative to the sequence $(\lambda_{k})$ of eigenvalues of problem (1.1), constructed via the Fadell–Rabinowitz cohomological index. See Theorem 3.3 for definition and properties of this sequence.
Theorem 1.3.
Assume $N\geq 3$, ($\mathcal{H}$). Suppose that for some $k,m\in\mathbb{N}$ it holds
$\lambda_{k}=\lambda_{k+1}=\dots=\lambda_{k+m-1}<\lambda_{k+m}.$
Consider $f$ as in (1.9), with $\mu\geq 0$, $\eta_{1}$ satisfying (1.10) and $(\eta_{2},r_{2})$ superscaled satisfying one of the conditions (1.11)-(1.13). Then, there is a number $\delta_{k}>0$ such that problem (1.4) possesses at least $m$ distinct pairs of nontrivial solutions with positive energy whenever $\lambda\in\bigl(\lambda_{k}-\delta_{k},\;\lambda_{k}\bigr).$
As an immediate consequence of this theorem we get the following
Corollary 1.4.
If $\lambda_{k}<\lambda_{k+1}$ for some $k\geq 1$, then, for $f$ as in Theorem 1.3, there is $\delta_{k}>0$ such that problem (1.4) has a pair of nontrivial solutions at a positive energy level for all $\lambda\in(\lambda_{k}-\delta_{k},\lambda_{k})$.
Instead of assuming that $\lambda$ is close to $\lambda_{k}$ in Theorem 1.3, one may alternatively require $\mu$ to be sufficiently large. In this case, however, the exponent $r_{2}$ is necessarily restricted to the subcritical range. We have the following result.
Theorem 1.5.
Assume $N\geq 3$ and ($\mathcal{H}$). Consider $f$ as in (1.9). Then, for every $\lambda\in\mathbb{R}$, $\eta_{1}$ satisfying (1.10) and $(\eta_{2},r_{2})$ superscaled satisfying one of the conditions (1.11)-(1.12), and $m\in\mathbb{N}$ there exists $\mu_{m}>0$ such that, for all $\mu>\mu_{m}$, problem (1.4) admits at least $m$ distinct pairs of nontrivial solutions with positive energy. In particular, the number of these solutions tends to infinity as $\mu\to\infty$.
From now on, let us turn our attention to nonlinearities that display at least a subscaled pair $(\eta,r)$ in terms of Definition 1. The first result in this setting concerns a single subscaled term and gives existence of infinitely many solutions on negative energy levels.
Theorem 1.6.
Assume ($\mathcal{H}$). For a subscaled pair $(\eta,r)$, with $b<\eta<2$, consider
$f(x,t)=\frac{|t|^{r-2}t}{|x|^{\eta}},\quad a.e.\,x\in\mathbb{R}^{N},\forall t\in\mathbb{R}.$
Then problem (1.4) has a sequence of solutions $(u_{k})$ at negative levels such that $\Phi(u_{k})\nearrow 0$.
For the next Theorem, recall the notation given in (1.7).
Theorem 1.7.
Assume ($\mathcal{H}$). Consider
$f(x,t)=\lambda\frac{|t|^{p-2}t}{|x|^{a}}+\frac{|t|^{r-2}t}{|x|^{\eta}}-\frac{|t|^{{r_{1}}-2}t}{|x|^{\eta_{1}}},\quad a.e.\,x\in\mathbb{R}^{N},\forall t\in\mathbb{R},$
where $(\eta,r)$ is subscaled. Suppose also that $(\eta_{1},r_{1})$ is an admissible pair satisfying
$\ell(\eta,r)<\ell(\eta_{1},r_{1}),\quad\mbox{for}\,\,\lambda\leq 0\quad\mbox{or}\quad(\eta_{1},r_{1})\text{ superscaled }\quad{for}\,\,\lambda>0.$
Then problem (1.4) has a sequence of solutions $(u_{k})$ at negative levels such that $\Phi(u_{k})\nearrow 0$.
The next theorem concerns asymptotically scaled nonlinearities. It requires that $\lambda$ is not an eigenvalue of (1.1); that is, $\lambda$ must not only differ from the sequence of eigenvalues obtained in Theorem 3.3, but also be such that (1.1) admits no nontrivial solution. From Theorem 3.3, it follows immediately that there are no eigenvalues in the interval $(-\infty,\lambda_{1})$, where $\lambda_{1}>0$ is defined in that theorem.
Theorem 1.8.
Assume ($\mathcal{H}$). Consider
$f(x,t)=\lambda\frac{|t|^{p-2}t}{|x|^{a}}+\frac{|t|^{r-2}t}{|x|^{\eta}},\quad a.e.\,x\in\mathbb{R}^{N},\forall t\in\mathbb{R},$
where $(\eta,r)$ is subscaled. Suppose that $\lambda$ is not an eigenvalue for (1.1). Then problem (1.4) has a sequence of solutions $(u_{k})$ at negative levels such that $\Phi(u_{k})\nearrow 0$.
Notice that this theorem includes Theorem 1.6 as the particular case $\lambda=0$. Indeed, if $\lambda\leq 0$, the associated functional is coercive, and the proof proceeds exactly as in Theorem 1.6. The genuinely new aspect of the present theorem concerns the case $\lambda>0$, where coercivity is lost and the verification of the $(PS)$ condition, as well as the geometric properties of the functional, requires a different argument.
The next result still gives infinitely many solutions in negative levels, but now with a critical nonlinearity with positive sign, perturbed with a subscaled nonlinearity with a small parameter.
Theorem 1.9.
Assume $N\geq 3$ and ($\mathcal{H}$). Consider
$f(x,t)=\mu\frac{|t|^{r-2}t}{|x|^{\eta}}+\frac{|t|^{2^{*}_{\eta_{1}}-2}t}{|x|^{\eta_{1}}},\quad a.e.\,x\in\mathbb{R}^{N},\forall t\in\mathbb{R},$
where $(\eta,r)$ is subscaled and $0\leq\eta_{1}<2$. The there exists a $\mu^{*}>0$ such that problem (1.4) has a sequence of solutions $(u_{k})$ at negative levels with $\Phi(u_{k})\nearrow 0$ for each $\mu\in(0,\mu^{*})$.
Our final theorem establishes a multiplicity result for a critical problem involving both scaled and subscaled nonlinearities. As in Theorem 1.3 and Corollary 1.4, its statement requires referring to the sequence of eigenvalues associated with problem (1.1), whose existence is ensured by Theorem 3.3.
Theorem 1.10.
Assume $N\geq 3$ and ($\mathcal{H}$). Consider
$f(x,t)=\lambda\frac{|t|^{p-2}t}{|x|^{a}}-\mu\frac{|t|^{r-2}t}{|x|^{\eta}}+\frac{|t|^{2^{*}_{\eta_{1}}-2}t}{|x|^{\eta_{1}}},\quad a.e.\,x\in\mathbb{R}^{N},\forall t\in\mathbb{R},$
where $(\eta,r)$ is subscaled and $\eta_{1}$ satisfies (1.10). Suppose that $\lambda>\lambda_{k}$. Then there exists $\mu^{*}>0$ such that, for every $\mu\in(0,\mu^{*})$, problem (1.4) admits $k$ pairs of nontrivial solutions at positive energy levels.
Organization of the paper
The structure of the remainder of the paper is as follows: In Section 2 we give some preliminary results such as the embedding theorems that justify Definition 1, and a Pohozaev Identity. Section 3 is devoted to the abstract framework of Mercuri and Perera [23]. We verify that our framework satisfy the structural hypotheses of the theory, and we construct the nonlinear eigenvalues $(\lambda_{k})$ of problem (1.1) using the Fadell–Rabinowitz cohomological index. Section 4 contains all proofs of the main results. We first establish general Palais–Smale and boundedness criteria. These preliminary tools are then used to treat, in a unified manner, some classes of nonlinearities considered in the paper: superscaled perturbations in the subcritical and critical cases; combinations of scaled, subscaled, and superscaled terms; and problems involving the nonlinear eigenvalues $(\lambda_{k})$. Section 5 is devoted to the radial setting. Using the improved radial Caffarelli–Kohn–Nirenberg inequalities, we obtain strictly larger embedding ranges for $E_{b,\mathrm{rad}}^{q}$ and extend the admissible regimes for $(\eta,r)$, particularly in the case $\eta<b$ and in dimension $N=2$, where no nonradial compact embeddings are available. We explain how all main existence and multiplicity theorems admit strengthened radial versions under these improved embedding intervals. Finally, Section 6 contains an appendix where we prove the density of smooth functions in the underlying functional space used throughout this paper.
2 Preliminaries
In this section we establish several basic ingredients needed for our analysis. We begin with the embedding results that enable the use of variational methods in the proof of our main theorems. We then derive a technical regularity lemma and a Pohozaev identity, which also play a crucial role in obtaining our results.
2.1 Embeddings in weighted Lebesgue spaces
Due to the Caffarelli - Kohn - Nirenberg inequality (see [7]), for fixed numbers
$r>0,\quad\theta\in[0,1],\quad q\geq 1,\quad\sigma\in\mathbb{R},\quad\gamma=\theta\sigma-(1-\theta)\frac{b}{q}$ (2.1)
there exists $C>0$ such that
$\||x|^{\gamma}u\|_{r}\leq C\|\nabla u\|^{\theta}_{2}\||x|^{\frac{-b}{q}}u\|_{q}^{1-\theta},\quad\forall u\in C^{\infty}_{c}(\mathbb{R}^{N})$ (2.2)
if, and only if, the following occur
$0<\frac{1}{r}+\frac{\gamma}{N}=\theta\frac{N-2}{2N}+(1-\theta)\frac{N-b}{qN},$ (2.3)
$\sigma\leq 0\,\,\,\text{if}\,\,\,\theta>0,\quad\text{and}\quad\sigma\geq-1\,\,\,\text{if}\,\,\,\theta>0\,\,\,\text{and}\,\,\,\frac{1}{r}+\frac{\gamma}{N}=\frac{N-2}{2N}.$ (2.4)
Moreover, $C$ is bounded if $\sigma\in[-1,0]$ and the other parameters are in a compact set.
We shall also make use of the following proposition, due to Mallick and Nguyen [22], which is stated below for convenience with the particular choice of parameters relevant to the present work.
Proposition 2.1.
[22, Proposition 24] Let $N\geq 1$, $q\geq 1$, $r\geq 1$, $0<\theta<1$, $0<b<N$ and $\gamma\in\mathbb{R}$. Define $\sigma$ by (2.1) and assume that $\sigma<0$, (2.3) holds and
$\frac{N-2}{2N}\neq\frac{N-b}{qN}.$
Suppose that the embedding $W^{1,2}(B_{1})\cap L^{q}(B_{1})$ into $L^{r}(B_{1})$ is compact. Let $(u_{n})_{n}\subset L^{1}_{\mathrm{loc}}(\mathbb{R}^{N})$ be a sequence of compact support functions such that $\bigl(\|\nabla u_{n}\|_{2}\bigr)$ and $\bigl(\|\,|x|^{-\frac{b}{q}}u_{n}\|_{q}\bigr)$ are bounded. Then $\bigl(|x|^{\gamma}u_{n}\bigr)$ has a convergent subsequence in $L^{r}(\mathbb{R}^{N})$.
The next result combines the classical CKN inequality with some new insights in [22] and brings some embedding results for the space $E^{q}_{b}$. Recal the definition of $2^{*}_{b,q,\eta}$ given in (1.5).
Lemma 2.2.
Suppose $N\geq 2$, let $0<b<2<q<2^{*}_{b}$ and $0\leq\eta<N$. For $N=2$, assume, additionally, that $b\leq\eta$. Fix $r>0$ such that
$r\in\left(2^{*}_{b,q,\eta},2^{*}_{\eta}\right)$
and $\theta\in(0,1)$ such that
$\frac{1}{r}=\theta\frac{N-2}{2(N-\eta)}+(1-\theta)\frac{N-b}{q(N-\eta)}.$ (2.5)
Moreover, if we assume $\eta\leq b$ we can consider $r=2^{*}_{b,q,\eta}$ and, if $\eta\leq 2$ and $N\geq 3$, the other endpoint $r=2^{*}_{\eta}$. Then, it holds
$\left(\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx\right)^{\frac{1}{r}}\leq C\left(\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx\right)^{\frac{\theta}{2}}\left(\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx\right)^{\frac{1-\theta}{q}},\quad\forall u\in E^{q}_{b}(\mathbb{R}^{N}).$ (2.6)
In particular, we have $E^{q}_{b}\hookrightarrow L_{\eta}^{r}(\mathbb{R}^{N})$ for every $r\geq 1$ in the above interval. Moreover, for $N=2$ with $\eta>b$, or for $N\geq 3$, these embeddings are compact except at the endpoints of the interval.
Proof.
Fix $r$ in the closure of one of the intervals and choose $\theta\in[0,1]$ so that (2.5) holds for this value of $r$. Consider $\gamma=-\eta/r$ so that (2.3) can be written as (2.5). Let $\sigma\in\mathbb{R}$ be given such that $\gamma=\theta\sigma-(1-\theta)b/q$, as in (2.1). Then, if $\theta=0$ we see that this expression of $\gamma$ and (2.5) imply
$\frac{\eta}{r}=-\gamma=\frac{b}{q}\quad\text{and}\quad\frac{1}{r}=\frac{N-b}{q(N-\eta)}\qquad\Longrightarrow\qquad\eta=b\,\,\,\text{and}\,\,\,r=q.$
If $\theta>0$, due to (2.5) we get
$\sigma=\frac{-\eta}{\theta r}+\frac{(1-\theta)b}{\theta q}=-\frac{\eta(N-2)}{2(N-\eta)}-\frac{(1-\theta)N(\eta-b)}{\theta q(N-\eta)}.$
We observe that for $N=2$, $\theta\in(0,1)$ and $0\leq\eta<N$ we get $sgn(\sigma)=sgn(b-\eta)$. This means that (2.4) only holds if $b\leq\eta$. Notice that for $\theta\in(0,1)$, we have $\sigma<0$ if $b<\eta$ and $\sigma=0$ if $b=\eta$. Then, for $\eta\geq b$, the CKN inequality (2.2) is valid with
$\frac{1}{r}=(1-\theta)\frac{N-b}{q(N-\eta)},$
where $\theta\in(0,1)$ if $\eta>b$ and $\theta\in[0,1)$ if $\eta=b$. On the other hand, for $N\geq 3$ we get $\sigma\leq 0$ for any $\theta\in(0,1]$ if $b\leq\eta$ while for $0\leq\eta<b$ it holds
$\sigma\leq 0\quad\Longleftrightarrow\quad\theta\geq\theta_{1}:=\frac{2N(b-\eta)}{\eta q(N-2)+2N(b-\eta)}\in(0,1].$
By placing $\theta=\theta_{1}$ in (2.5) we obtain $r=2^{*}_{b,q,\eta}$. Moreover, when $\theta=1$ and $\eta\leq 2$ we get $\sigma\geq-1$. So, $\sigma$ satisfies (2.4) and then (2.2) occurs for $r$ between $2^{*}_{b,q,\eta}$ and ${2(N-\eta)}/{(N-2)}$, including ${2(N-\eta)}/{(N-2)}$ only in case $\eta\leq 2$, and including $2^{*}_{b,q,\eta}$} only in case $\eta\leq b$.
Under these conditions, we conclude that (2.6) follows by density.
It remains to prove that these embeddings are compact (except at the endpoints). Let $(u_{n})$ be a bounded sequence of functions in $E^{q}_{b}$ with $u_{n}\in C_{c}^{\infty}(\mathbb{R}^{N})$. This means that the sequences $(\|\nabla u_{n}\|_{2})$ and $\bigl(\|\,|x|^{-b/q}u_{n}\|_{q}\bigr)$ are bounded. Then, if $N=2$ with $\eta>b$ or $N\geq 3$, fixing $r\geq 1$ in one of those (open) intervals and setting $\gamma=-\eta/r$, all the chosen parameters satisfy the assumptions of Proposition 2.1. Notice that the space $W^{1,2}(B_{1})\cap L^{q}(B_{1})$ is compactly embedded into $L^{k}(B_{1})$ for every $1\leq k<\max\{q,\,2N/(N-2)\}$ (or $k\geq 1$ if $N=2$), and $r$ lies within this range. Hence, $\bigl(|x|^{-\eta/r}u_{n}\bigr)$ admits a convergent subsequence in $L^{r}(\mathbb{R}^{N})$, which completes the proof. ∎
It is worth noting that the upper endpoint $r=2^{*}_{\eta}$ of the embedding ranges obtained in Lemma 2.2 cannot be improved. Indeed, the standard concentration argument applies: fix a nontrivial radial function supported in an annulus and consider the rescalings $u_{\lambda}(x)=\lambda^{(N-2)/2}u(\lambda x)$. This family is bounded in $E_{b}^{q}$, but no subsequence can converge in $L^{2^{*}_{\eta}}_{\eta}(\mathbb{R}^{N})$. Thus compactness fails precisely at the critical exponent $2^{*}_{\eta}$, as in the classical Sobolev setting. In turn, this failure of compactness shows that $2^{*}_{\eta}$ is also the optimal endpoint for the continuous embedding, as expected.
In contrast, the lower threshold exponent $2^{*}_{b,q,\eta}$, which marks the beginning of the admissible regime, is not known to be optimal. The scaling construction used above does not give non-compactness at this value, and whether continuity or compactness extends up to this lower endpoint appears to remain an open question.
The corollary below shows the compactness of the embeddings of the underlying functional spaces associated with the nonlinear eingenvalue problem (1.1). This property is crucial in the development of the eigenvalue theory based on the $\mathbb{Z}_{2}$-cohomological index of Fadell and Rabinowitz [12].
Corollary 2.3.
Assume ($\mathcal{H}$). Then we have $E^{q}_{b}(\mathbb{R}^{N})\subset L^{p}_{a}(\mathbb{R}^{N})$ with compact embedding, for $a$ as defined in (1.3).
Proof.
For $N\geq 2$ we verify that
$2^{*}_{b,q,a}<p\quad\Longleftrightarrow\quad q[(N-2)(q-2)-2(2-b)]<p[(N-2)(q-2)-2(2-b)]$
and for $N\geq 3$ we get
$p<2^{*}_{a}\quad\Longleftrightarrow\quad p[(N-2)(q-2)-2(2-b)]<2[(N-2)(q-2)-2(2-b)].$
Thus, since $p\in(2,q)$ and $q<2^{*}_{b}$, we obtain $2^{*}_{b,q,a}<p<2^{*}_{a}.$
2.2 Pohozaev Identity and nonexistence results
In this section we establish a Pohozaev identity that plays a crucial role in our results. The proof follows the ideas of [28]. This identity also yields conditions ensuring nonexistence of solutions to problem (1.4).
Proposition 2.4.
Suppose $N\geq 2$ and let $0<b<2<q<2^{*}_{b}$. Let $u\in E^{q}_{b}$ be a weak solution for (1.4) where $f:\mathbb{R}^{N}\times\mathbb{R}\to\mathbb{R}$ is a Carathéodory function given by
$f(x,t)=\sum_{j=1}^{m}c_{j}\frac{|t|^{r_{j}-2}t}{|x|^{\eta_{j}}}$
for some $\eta_{j}\in[0,2)$ and $1<r_{j}\leq 2^{*}_{\eta_{j}}$ ($r_{j}\in(1,\infty)$ if $N=2$). Then $u\in C^{2}(\mathbb{R}^{N}\backslash\{0\})$ and satisfies the Pohozaev type identity
$\frac{N-2}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\frac{N-b}{q}\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx=\sum_{j=1}^{m}\frac{c_{j}(N-\eta_{j})}{r_{j}}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{j}}}{|x|^{\eta_{j}}}dx.$ (2.7)
Proof.
Let $u\in E_{b}^{q}$ be a weak solution of (1.4). We first show that $u\in C^{2}(\mathbb{R}^{N}\setminus\{0\})$. Denote $h(x,u)=f(x,u)-|x|^{-b}{|u|^{q-2}u}.$ Notice that we have $h(x,u)\in L^{r}_{\mathrm{loc}}(\mathbb{R}^{N}\setminus\{0\})$ for every $r\geq 1$. When $N=2$, this follows directly from the embeddings in Lemma 2.2; when $N\geq 3$, it follows from Lemma B.3 in [31]. Standard Calderón–Zygmund estimates then give $u\in W^{2,r}_{\mathrm{loc}}(\mathbb{R}^{N}\setminus\{0\})$ for all $r\geq 1$, and Schauder estimates yield $u\in C^{2}(\mathbb{R}^{N}\setminus\{0\}).$
To prove the Pohozaev identity, we fix $0<r<R$ and set $A_{r,R}=\{x\in\mathbb{R}^{N}:r<|x|<R\}$. Multiplying the equation (1.4) by $x\cdot\nabla u$ and integrating over $A_{r,R}$, an application of the Green’s identity gives
$\frac{N-2}{2}\int_{A_{r,R}}|\nabla u|^{2}dx+\frac{N-b}{q}\int_{A_{r,R}}\frac{|u|^{q}}{|x|^{b}}dx=\sum_{j=1}^{m}c_{j}\,\frac{N-\eta_{j}}{r_{j}}\int_{A_{r,R}}\frac{|u|^{r_{j}}}{|x|^{\eta_{j}}}dx+\beta(R)-\beta(r),$
where
$\beta(t)=\frac{t}{2}\int_{\partial B_{t}}|\nabla u|^{2}dS-t\int_{\partial B_{t}}|\nabla u\cdot\nu|^{2}dS+\frac{t}{q}\int_{\partial B_{t}}\frac{|u|^{q}}{|x|^{b}}dS-\sum_{j=1}^{m}c_{j}\frac{t}{r_{j}}\int_{\partial B_{t}}\frac{|u|^{r_{j}}}{|x|^{\eta_{j}}}dS.$
Since all integrands belong to $L^{1}(\mathbb{R}^{N})$, we may choose sequences $r_{n}\to 0^{+}$ and $R_{n}\to\infty$ such that $\beta(r_{n})\to 0$ and $\beta(R_{n})\to 0$. Letting $n\to\infty$ yields the identity (2.7). ∎
This Pohozaev identity leads to the following nonexistence result, which is similar to [30, Theorem 9] (see also [26, Theorem 4.11]).
Proposition 2.5.
Assume $0<b<2<q<2^{*}_{b}$ and consider $\eta\in[0,2)$. Let $u\in E^{q}_{b}\cap L^{r}_{\eta}(\mathbb{R}^{N})$ be a solution of the problem
$\Delta u+\frac{|u|^{q-2}u}{|x|^{b}}=\frac{|u|^{r-2}u}{|x|^{\eta}}\qquad\text{in }\mathbb{R}^{N}.$ (2.8)
Assume that $u\in L^{\infty}_{loc}(\mathbb{R}^{N}\backslash\{0\})$ if $r>2^{*}$. Then $u\equiv 0$ if one of the following conditions occur:
  1. 1.
    $r\geq 2^{*}_{\eta}$ or $1<r\leq q(N-\eta)/(N-b)$, if $N\geq 3$;
  2. 2.
    $1<r\leq q(N-\eta)/(N-b)$, if $N=2$.
Proof.
As in the proof of Proposition 2.4 we see that $u\in C^{2}(\mathbb{R}^{N}\backslash\{0\})$ if $r\leq 2^{*}$. If $r>2^{*}$ and we assume $u\in L^{\infty}_{loc}(\mathbb{R}^{N}\backslash\{0\})$ then we reach the same regularity. So, the Pohozaev identity (2.7) holds for $u$, which means that
$\frac{N-2}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\frac{N-b}{q}\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx=\frac{N-\eta}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx.$
On the other hand, from (2.8) we get
$\frac{N-\eta}{r}\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\frac{N-\eta}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx=\frac{N-\eta}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx.$
Subtracting one equation from another we obtain
$\left(\frac{N-2}{2}-\frac{N-\eta}{r}\right)\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\left(\frac{N-b}{q}-\frac{N-\eta}{r}\right)\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx=0.$
Therefore, from assumptions 1. or 2. it follows that $u=0$. ∎
Notice that Lemma 2.2 and Proposition 2.5 show that $2_{\eta}^{*}$ is a critical threshold for the embeddings and also that it plays the same role for the existence of nontrivial solutions, which is the expected behavior. Even more interesting is the fact that the number $q(N-\eta)/(N-b)$ also emerges as a threshold for existence. Observe that when $\eta\geq b$ we have that the critical lower endpoint for the embeddings in Lemma 2.2 coincides with the corresponding nonexistence threshold. In contrast, when $\eta<b$ there is a gap
$\frac{q(N-\eta)}{N-b}<r<2^{*}_{b,q,\eta}.$
This gap is shortened in the radial case (see Section 5).
3 Scaled operators on $E^{q}_{b}(\mathbb{R}^{N})$
Consider the definition of a scaling on a Banach space, as given in [23, Definition 1.1].
Definition 2.
Let $W$ be a reflexive Banach space. A scaling on $W$ is a continuous mapping
$W\times[0,\infty)\to W,\quad(u,t)\mapsto u_{t}$
satisfying
  1. ($H_{1}$)
    $(u_{t_{1}})_{t_{2}}=u_{t_{1}t_{2}}$ for all $u\in W$ and $t_{1},t_{2}\geq 0$,
  2. ($H_{2}$)
    $(\tau u)_{t}=\tau u_{t}$ for all $u\in W$, $\tau\in\mathbb{R}$ and $t\geq 0$,
  3. ($H_{3}$)
    $u_{0}=0$ and $u_{1}=u$ for all $u\in W$,
  4. ($H_{4}$)
    $u_{t}$ is bounded on bounded sets in $W\times[0,\infty)$,
  5. ($H_{5}$)
    $\exists\ell>0$ such that $\|u_{t}\|=O(t^{\,\ell})$ as $t\to\infty$, uniformly in $u$ on bounded sets.
We have the following result.
Lemma 3.1.
Suppose $N\geq 2$ and let $0<b<2<q<2^{*}_{b}$. Consider $u_{t}$ as defined in (1.2). The map $E^{q}_{b}\times[0,\infty)\to E^{q}_{b}$ given by $(u,t)\mapsto u_{t}$ is a scaling on $E^{q}_{b}$.
Proof.
It is straightforward to verify that ($H_{1}$)-($H_{3}$) hold. To prove ($H_{4}$) and ($H_{5}$), recall the notation given in (1.7). Then
$\|\nabla u_{t}\|_{2}^{2}=t^{\,\ell}\|\nabla u\|_{2}^{2}\quad\text{and}\quad\int_{\mathbb{R}^{N}}\frac{|u_{t}|^{q}}{|x|^{b}}dx=t^{\,\ell}\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx$
for all $u\in E^{q}_{b}$ and $t\geq 0$. Therefore
$\|u_{t}\|\leq\max\{t^{\,\ell/2},t^{\,\ell/q}\}\|u\|,\quad\forall(u,t)\in E^{q}_{b}\times[0,\infty),$ (3.1)
which implies ($H_{4}$) and ($H_{5}$) are satisfied with $\ell=(q\delta+b)-N$. Continuity follows then as in [23, Lemma 3.1] as a result of the density of $C^{\infty}_{c}(\mathbb{R}^{N})$ in $E^{q}_{b}$. ∎
Let $\tilde{h}\in C(E^{q}_{b},(E^{q}_{b})^{*})$. We say that $\tilde{h}$ is a potential operator if there exists a functional $\tilde{H}\in C^{1}(E^{q}_{b},\mathbb{R})$, called a potential of $\tilde{h}$, such that $\tilde{h}=\tilde{H}^{\prime}$. Without loss of generality, we may take $\tilde{H}(0)=0$ (if necessary, by considering $\tilde{H}-\tilde{H}(0)$).
An odd potential operator $\tilde{h}$ is said to be a scaled operator provided it maps bounded subsets of $E^{q}_{b}$ into bounded subsets of $(E^{q}_{b})^{*}$ and, for some exponent $\ell\in\mathbb{R}$, satisfies
$\tilde{h}(u_{t})v_{t}=t^{\,\ell}\tilde{h}(u)v\quad\text{for all }u,v\in E^{q}_{b}\text{ and }t\geq 0.$ (3.2)
According to [23, Proposition 2.2], whenever $\tilde{h}$ is a potential operator, the corresponding potential $\tilde{H}$ with $\tilde{H}(0)=0$ can be expressed as
$\tilde{H}(u)=\int_{0}^{1}\tilde{h}(\tau u)\,u\,d\tau,\qquad u\in E^{q}_{b}.$
It follows from this representation that $\tilde{H}$ is even whenever $\tilde{h}$ is odd, and that $\tilde{H}$ maps bounded sets to bounded sets if the same holds for $\tilde{h}$. In addition, when $\tilde{h}$ satisfies the scaling condition (3.2), the potential $\tilde{H}$ inherits the same homogeneity property:
$\tilde{H}(u_{t})=t^{\,\ell}\tilde{H}(u),\qquad u\in E^{q}_{b},\,t\geq 0.$
With the above notions at hand, we now introduce the odd potential operators
$\mathcal{A},\,\mathcal{B}:E^{q}_{b}\to(E^{q}_{b})^{*},$
defined respectively by
$\mathcal{A}(u):=-\Delta u+\frac{|u|^{q-2}u}{|x|^{b}}\quad\text{and}\quad\mathcal{B}(u):=\frac{|u|^{p-2}u}{|x|^{a}}.$ (3.3)
That means
$\mathcal{A}(u)v=\int_{\mathbb{R}^{N}}\nabla u\nabla v\,dx+\int_{\mathbb{R}^{N}}\frac{|u|^{q-2}u\,v}{|x|^{b}}\,dx,$
and
$\mathcal{B}(u)v=\int_{\mathbb{R}^{N}}\frac{|u|^{p-2}u\,v}{|x|^{a}}\,dx.$
Here $2<p<q$ and the parameters satisfy $0<b<a<2$, as specified in (1.3). Both $\mathcal{A}$ and $\mathcal{B}$ fulfill the scaling condition (3.2) with the same exponent $\ell$ as defined in (1.7), which is consistent with assumption ($H_{5}$). Furthermore, the continuous embedding
$E^{q}_{b}\hookrightarrow L^{p}_{a}(\mathbb{R}^{N}),$
provided by Corollary 2.3, ensures that $\mathcal{B}$ maps bounded sets in $E^{q}_{b}$ to bounded sets in $(E^{q}_{b})^{*}$. The same property is immediate for $\mathcal{A}$. Consequently, both $\mathcal{A}$ and $\mathcal{B}$ are scaled potential operators.
Notice that if $u\in E^{q}_{b}$ satisfies $\mathcal{A}(u)=\lambda\,\mathcal{B}(u)$, that is, if $u$ is a solution of problem (1.1), then, by the scaling properties of $\mathcal{A}$ and $\mathcal{B}$, its rescaled function $u_{t}$ is also a solution for every $t\geq 0$. Associated with the operators $\mathcal{A}$ and $\mathcal{B}$, we introduce the corresponding potentials
$I,\,J:E^{q}_{b}\longrightarrow\mathbb{R},$
defined by
$I(u)=\frac{1}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}\,dx+\frac{1}{q}\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}\,dx,\qquad J(u)=\frac{1}{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}\,dx.$ (3.4)
In order to apply the abstract critical point framework developed in [23], we must verify that the operators $\mathcal{A}$ and $\mathcal{B}$, as well as their corresponding potentials $I$ and $J$, satisfy the structural properties required in that setting. They are the following:
  1. ($H_{6}$)
    $\mathcal{A}(u)u>0$ for all $u\in E^{q}_{b}\backslash\{0\}.$
  2. ($H_{7}$)
    If $u_{n}\rightharpoonup u$ in $E$ and $\mathcal{A}(u_{n})(u_{n}-u)\to 0$ then $(u_{n})$ has a subsequence that converges strongly to $u$.
  3. ($H_{8}$)
    $\mathcal{B}(u)u>0$ for all $u\in E^{q}_{b}\backslash\{0\}.$
  4. ($H_{9}$)
    If $u_{n}\rightharpoonup u$ in $E^{q}_{b}$ then $\mathcal{B}(u_{n})\to\mathcal{B}(u)$ in $(E^{q}_{b})^{*}$.
  5. ($H_{10}$)
    II is coercive, i.e., $I(u)\to\infty$ as $\|u\|\to\infty.$
  6. ($H_{11}$)
    For each $u\in E^{q}_{b}\backslash\{0\}$ there is an unique $t>0$ such that $I(tu)=1.$
  7. ($H_{12}$)
    Every solution $u$ of equation (1.1) satisfies $I(u)=\lambda J(u).$
We check these conditions in the following lemma.
Lemma 3.2.
Assume ($\mathcal{H}$). Let $\mathcal{A},\mathcal{B}$ and $I$, $J$ as defined in (3.3) and (3.4). Then, the conditions ($H_{6}$)-($H_{12}$) hold true.
Proof.
[…]
$\mathcal{A}(u_{n})(u_{n}-u)=\int_{\mathbb{R}^{N}}\nabla u_{n}\nabla(u_{n}-u)dx+\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{q-2}u_{n}(u_{n}-u)}{|x|^{b}}\,dx.$
Since $u_{n}\rightharpoonup u$ in $E^{q}_{b}$, we also have $\nabla u_{n}\rightharpoonup\nabla u$ weakly in $L^{2}(\mathbb{R}^{N})$. This yields
$\int_{\mathbb{R}^{N}}\nabla u_{n}\nabla(u_{n}-u)dx=\|\nabla(u_{n}-u)\|_{2}^{2}+o_{n}(1).$
Moreover, since the embedding $E^{q}_{b}\hookrightarrow L^{q}_{b}(\mathbb{R}^{N})$ is continuous (Lemma 2.2), we must have $u_{n}\rightharpoonup u$ in $L^{q}_{b}(\mathbb{R}^{N})$ as well. So, a standard duality argument gives that
$\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{q-2}u_{n}u}{|x|^{b}}\,dx\rightarrow\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}\,dx,$
up to a subsequence. Then, the Brezis-Lieb lemma (with weights) gives
$\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{q-2}u_{n}(u_{n}-u)}{|x|^{b}}\,dx=\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{q}}{|x|^{b}}dx+o_{n}(1).$
Thus, up to a subsequence, we have
$\mathcal{A}(u_{n})(u_{n}-u)=\|\nabla(u_{n}-u)\|_{2}^{2}+\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{q}}{|x|^{b}}dx+o_{n}(1),$
which concludes the proof of ($H_{7}$).
Finally, suppose that $u$ is a nontrivial solution of (1.1). Then, $\mathcal{A}(u)=\lambda\mathcal{B}(u)$ (notice that, due to ($H_{6}$) and ($H_{8}$), we have $\lambda>0$). Then, using $u$ as a test function we have
$\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx=\lambda\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx.$ (3.5)
Moreover, the Pohozaev Identity (2.7) gives
$\frac{N-2}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\frac{N-b}{q}\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx=\frac{\lambda(N-a)}{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx.$ (3.6)
Multiply (3.6) by q2q-2 and (3.5) by 2b2-b and subtract one from another. Then, divide the resulting equation by 2(Nb)q(N2)2(N-b)-q(N-2), which is nonzero, and we arrive at ($H_{12}$). ∎
3.1 Scaled eigenvalue problem
We continue to explore the scaled eigenvalue problem $\mathcal{A}=\lambda\mathcal{B}$, i.e. problem (1.1) with $a>0$ as defined in (1.3). Whenever (1.1) admits a nontrivial solution $u\in E^{q}_{b}$, we refer to the corresponding parameter $\lambda$ as an eigenvalue, and to $u$ as an eigenfunction associated with $\lambda$. It was shown above that the hypotheses ($H_{6}$)–($H_{12}$) are fulfilled by the scaled operators $\mathcal{A}$ and $\mathcal{B}$, together with their respective potentials $I$ and $J$, as given in (3.4). We denote by $\sigma(\mathcal{A},\mathcal{B})$ the set of all eigenvalues of problem (1.1), and call it the spectrum of the pair of scaled operators $(\mathcal{A},\mathcal{B})$. From ($H_{6}$) and ($H_{8}$), it follows that $\sigma(\mathcal{A},\mathcal{B})\subset(0,\infty).$ We set
$\mathcal{M}=\{u\in E^{q}_{b}:I(u)=1\},$
[…]
$\pi(u)=u_{t_{u}},\quad t_{u}=\left({I(u)}\right)^{-1/\ell}$ (3.7)
where $u_{t}$ is the scaling defined in (1.2) and $\ell$ is given in (1.7). We also define
$\Psi(u)=\frac{1}{J(u)},\quad u\in E^{q}_{b}\backslash\{0\},\quad\mbox{and}\quad\tilde{\Psi}=\Psi_{|_{\mathcal{M}}}.$ (3.8)
We observe that $\mathcal{M}$ is a complete, symmetric, and bounded $C^{1}$-Finsler manifold, and that the eigenvalues of problem (1.1) coincide with the critical values of $\tilde{\Psi}$, as established in [23, Proposition 2.5]. Let $\mathcal{F}$ denote the family of symmetric subsets of $\mathcal{M}$, and let $i(Y)$ represent the cohomological index of $Y\in\mathcal{F}$ in the sense of Fadell and Rabinowitz (see [12]). For each $k\geq 1$, define
$\mathcal{F}_{k}=\{\,Y\in\mathcal{F}:i(Y)\geq k\,\},$
and set
$\lambda_{k}:=\inf_{Y\in\mathcal{F}_{k}}\sup_{u\in Y}\Psi(u).$
The following theorem follows directly from an abstract variational result proved in [23]. It guarantees that the sequence $(\lambda_{k})$ consists of nonlinear eigenvalues of (1.1) and provides additional properties, which are stated below.
Theorem 3.3 ([23], Theorem 1.3).
Assume ($\mathcal{H}$). Then $\lambda_{k}\nearrow\infty$ is a sequence of eigenvalues of (1.1). Moreover,
  1. (i)
    The first eigenvalue is given by
    $\lambda_{1}=\min_{u\in\mathcal{M}}\tilde{\Psi}(u)>0.$
  2. (ii)
    If $\lambda_{k}=\cdots=\lambda_{k+m-1}=\lambda$, then $i(E_{\lambda})\geq m$, where $E_{\lambda}$ is the set of eigenfunctions associated with $\lambda$ that lie on $\mathcal{M}$.
  3. (iii)
    If $\lambda_{k}<\lambda<\lambda_{k+1}$, then
    $i(\tilde{\Psi}^{\lambda_{k}})=i(\mathcal{M}\backslash\tilde{\Psi}_{\lambda})=i(\tilde{\Psi}^{\lambda})=i(\mathcal{M}\backslash\tilde{\Psi}_{\lambda_{k+1}})=k,$
    where $\tilde{\Psi}^{\alpha}=\{u\in\mathcal{M}:\tilde{\Psi}(u)\leq\alpha\}$ and $\tilde{\Psi}_{\alpha}=\{u\in\mathcal{M}:\tilde{\Psi}(u)\geq\alpha\}$ for $\alpha\in\mathbb{R}$.
4 Proof of the main theorems
Our existence and multiplicity results for problem (1.4) will be obtained by means of critical point theorems, primarily those in [23] together with an abstract result from [25]. For the reader’s convenience, several of these tools are recalled below (see Propositions 4.6, 4.8 and Corollary 4.7).
As is standard in critical point theory, compactness assumptions such as the Palais–Smale condition, either in its global or localized form, are required for the application of these results. We begin by establishing that such compactness properties indeed hold for the functionals considered in this work. We then prove the main theorems, organizing the presentation according to the growth regimes involved: subscaled, superscaled, and subcritical or critical, as well as combinations of these settings.
4.1 On the Palais-Smale condition
This subsection is devoted to verifying the Palais-Smale condition for the functionals defined in (1.6). We start with a subcritical superscaled case. See Definition 1 for the definition of an admissible and superscaled pair $(\eta,r)$.
Lemma 4.1.
Under the assumptions of Theorem 1.1, the functional $\Phi$ defined in (1.6) satisfies the (PS) condition.
Proof.
Here we have $G(x,t)=|x|^{-\eta}H(t)$ in the definition of $\Phi$. Let $c\in\mathbb{R}$ and $(u_{n})$ in $E^{q}_{b}$ a $(PS)_{c}$ sequence for $\Phi,$ namely
$\Phi(u_{n})=c+o(1)\quad\mbox{and}\quad\|\Phi^{\prime}(u_{n})\|_{(E^{q}_{b})^{*}}=o(1),\quad\mbox{as}\,\,n\to\infty.$
To show that it contains a strongly convergent subsequence, we first show that it is bounded in $E^{q}_{b}$. In fact, suppose, arguing by contradiction, that $\|u_{n}\|\to\infty$ for a subsequence. We use the notation (3.7) and set
$t_{n}=t_{u_{n}}=[I(u_{n})]^{-1/\ell},\quad\tilde{u}_{n}=\pi(u_{n})=(u_{n})_{t_{n}}\quad\mbox{and}\quad\tilde{t}_{n}=t_{n}^{-1}=[I(u_{n})]^{1/\ell}.$
Since $I$ is coercive by ($H_{10}$), we have $\tilde{t}_{n}\to\infty$. Once $\mathcal{M}$ is bounded, so it is $(\tilde{u}_{n})$. Then, up to a subsequence, $\tilde{u}_{n}\rightharpoonup v$ for some $v\in E^{q}_{b}$. Since $(\tilde{u}_{n})_{\tilde{t}_{n}}=u_{n}$ by ($H_{1}$), and using (3.1) we may estimate
$\|u_{n}\|=\|(\tilde{u}_{n})_{\tilde{t}_{n}}\|=O((\tilde{t}_{n})^{\ell/2})\quad\text{as}\quad n\to\infty.$ (4.1)
Since $\Phi(u_{n})=c+o(1)$, we observe that
$\Phi_{\lambda}(u_{n})=\int_{\mathbb{R}^{N}}G(x,u_{n})\,dx+c+o(1).$
Therefore, recalling that
$\Phi_{\lambda}(u_{t})=I(u_{t})-\lambda J(u_{t})=t^{\,\ell}\Phi_{\lambda}(u),\quad\forall u\in E,\forall t\geq 0,$
we have $\Phi_{\lambda}(u_{n})=(\tilde{t}_{n})^{\ell}\Phi_{\lambda}(\tilde{u}_{n}).$ So, we can write
$\Phi_{\lambda}(\tilde{u}_{n})=(\tilde{t}_{n})^{-\ell}\int_{\mathbb{R}^{N}}G(x,u_{n})\,dx+o(1),$ (4.2)
and using (1.7) and (1.8) we obtain
$\Phi_{\lambda}(\tilde{u}_{n})\geq\frac{c(\tilde{t}_{n})^{-\ell}}{r}\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{r}}{|x|^{\eta}}\,dx+o(1)=\frac{c(\tilde{t}_{n})^{\ell(\eta,r)-\ell}}{r}\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{r}}{|x|^{\eta}}\,dx+o(1).$
Since $r>q$, it is readily seen that $(\eta,r)$ is superscaled, that is, $\ell(\eta,r)>\ell$. Then, the boundedness of $(\Phi(\tilde{u}_{n}))$ and the unboundedness of $(\tilde{t}_{n})$ imply that $\tilde{u}_{n}\to 0$ in $L^{r}_{\eta}(\mathbb{R}^{N})$. Consequently, $\tilde{u}_{n}\rightharpoonup 0$ in $E^{q}_{b}$. From the compactness of the embedding $E^{q}_{b}\hookrightarrow L^{p}_{a}(\mathbb{R}^{N})$, see Corollary 2.3, we have $\tilde{u}_{n}\to 0$ in $L^{p}_{a}(\mathbb{R}^{N})$. On the other hand, using
$\Phi_{\lambda}^{\prime}(u_{t})u_{t}=\mathcal{A}(u_{t})u_{t}-\lambda\mathcal{B}(u_{t})u_{t}=t^{\,\ell}\Phi_{\lambda}^{\prime}(u)u,\quad\forall u\in E,\forall t\geq 0,$
$\Phi^{\prime}(u_{n})u_{n}=o(\|u_{n}\|)$ and (4.1), we get
$\Phi^{\prime}_{\lambda}(\tilde{u}_{n})\tilde{u}_{n}={(\tilde{t}_{n})^{-\ell}}\int_{\mathbb{R}^{N}}g(x,u_{n})u_{n}dx+o((\tilde{t}_{n})^{\ell/2}).$ (4.3)
Finally, multiplying (4.2) by $r$ and subtracting (4.3), due to (1.8) we conclude that
$\left(\frac{r}{2}-1\right)$ $\int_{\mathbb{R}^{N}}|\nabla\tilde{u}_{n}|^{2}dx+\left(\frac{r}{q}-1\right)\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{q}}{|x|^{b}}dx$
$=\lambda\left(\frac{r}{p}-1\right)\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{p}}{|x|^{a}}dx+(\tilde{t}_{n})^{-\ell}\int_{\mathbb{R}^{N}}\left[rG(x,u_{n})-g(x,u_{n})u_{n}\right]dx+o(1)\leq o(1).$
This implies that $\tilde{u}_{n}\to 0$ in $E^{q}_{b}$, which is an absurd for $\tilde{u}_{n}\in\mathcal{M}$. Therefore, $({u}_{n})$ must be bounded in $E^{q}_{b}$. Now, we consider the potential operator $\tilde{f}:E^{b}_{q}\to(E^{b}_{q})^{*}$, induced by $f$, defined by
$\tilde{f}(u)v=\lambda\int_{\mathbb{R}^{N}}\frac{|u|^{p-2}uv}{|x|^{a}}dx+\int_{\mathbb{R}^{N}}g(x,u)vdx.$
Since $E^{q}_{b}$ is compactly embedded in $L^{p}_{a}(\mathbb{R}^{N})$, $L^{r_{1}}_{\eta}(\mathbb{R}^{N})$ and $L^{r_{2}}_{\eta}(\mathbb{R}^{N})$ for $r_{1},r_{2}$ as in (1.8), see Lemma 2.6, $\tilde{f}$ is a compact operator. If $(u_{n})$ is a bounded $(PS)_{c}$ sequence for $\Phi$, we may use [23, Proposition 2.3] by which $(u_{n})$ has a convergent subsequence, and this concludes the proof. ∎
Now we focus our attention on critical nonlinearities given by sums of powers of $u$ with singular weights, as in (1.9). For convenience, we write it here again:
$f(x,u)=\lambda\frac{|u|^{p-2}u}{|x|^{a}}+\frac{|u|^{2^{*}_{\eta_{1}}-2}u}{|x|^{\eta_{1}}}+\mu\frac{|u|^{r_{2}-2}u}{|x|^{\eta_{2}}},\quad\lambda,\mu\in\mathbb{R}.$ (4.4)
This means that $G(x,u)=\frac{1}{2^{*}_{\eta_{1}}}\frac{|u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}+\frac{1}{r_{2}}\frac{|u|^{r_{2}}}{|x|^{\eta_{2}}}$ on the definition of functional $\Phi$ in (1.6). Here, the pair $(\eta_{1},2^{*}_{\eta_{1}})$ brings a critical superscaled configuration to problem (1.4), whereas $(\eta_{2},r_{2})$ may be treated in either the subscaled or superscaled regime, even critical. See Definition 1.
Lemma 4.2.
Assume $N\geq 3$ and ($\mathcal{H}$). For $f$ given by (4.4), suppose that $\lambda\in\mathbb{R}$, $0\leq\eta_{1}<N-q(N-2)/2$ and $\mu\in\mathbb{R}$. Moreover suppose that the admissible pair $(\eta_{2},r_{2})$ and $\mu$ satisfy one of the following assumptions:
i)
$\mu\in\mathbb{R}$ and $(\eta_{2},r_{2})$ is scaled or subscaled, with $b<\eta_{2}$ ($r_{2}<2$ if $\eta_{2}=2$);
ii)
$\mu>0$, $0\leq\eta_{2}<\eta_{1}$ and $2^{*}_{\eta_{1}}\leq r_{2}\leq 2^{*}_{\eta_{2}}$;
iii)
$\mu>0$, $(\eta_{2},r_{2})$ is superscaled, $\eta_{2}\in[\eta_{1},2)$ (equality only if $\eta_{1}=\eta_{2}>b$) and $r_{2}<2^{*}_{\eta_{2}}$. Additionally, assume that
$r_{2}>q+\frac{(2^{*}_{\eta_{1}}-q)(b-\eta_{2})}{b-\eta_{1}}\quad\text{if}\quad\eta_{2}<b.$
Then every $(PS)$ sequence of the functional $\Phi$ defined in (1.6) is bounded in $E_{b}^{q}$.
Proof.
Let $c\in\mathbb{R}$ and $(u_{n})$ in $E^{q}_{b}$ a $(PS)_{c}$ sequence for $\Phi.$ To show that $(u_{n})$ is bounded in $E^{q}_{b}$ we start arguing as in Lemma 4.1. We assume, by contradiction, that $\|u_{n}\|\to\infty$ for a subsequence, and consider
$t_{n}=t_{u_{n}}=[I(u_{n})]^{-1/\ell},\quad\tilde{u}_{n}=\pi(u_{n})=(u_{n})_{t_{n}}\quad\mbox{and}\quad\tilde{t}_{n}=t_{n}^{-1}=[I(u_{n})]^{1/\ell}.$
We have $\tilde{t}_{n}\to\infty$, $\tilde{u}_{n}\in\mathcal{M}$ and, up to a subsequence, $\tilde{u}_{n}\rightharpoonup v$ for some $v\in E^{q}_{b}$. Since $(\tilde{u}_{n})_{\tilde{t}_{n}}=u_{n}$, as in (4.2) we obtain
$\Phi_{\lambda}(\tilde{u}_{n})=\frac{(\tilde{t}_{n})^{\ell_{1}-\ell}}{2^{*}_{\eta_{1}}}\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx+\frac{\mu(\tilde{t}_{n})^{\ell_{2}-\ell}}{r_{2}}\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{r_{2}}}{|x|^{\eta_{2}}}dx+o(1),$ (4.5)
where $\ell_{1}=\ell(\eta_{1},2^{*}_{\eta_{1}})$ and $\ell_{2}=\ell(\eta_{2},r_{2})$ (see (1.7)). Also, similar to (4.3), we get
$\Phi^{\prime}_{\lambda}(\tilde{u}_{n})\tilde{u}_{n}={(\tilde{t}_{n})^{\ell_{1}-\ell}}\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx+{\mu(\tilde{t}_{n})^{\ell_{2}-\ell}}\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{r_{2}}}{|x|^{\eta_{2}}}dx+o((\tilde{t}_{n})^{\ell/2}).$ (4.6)
Finally, multiplying (4.5) by $2^{*}_{\eta_{1}}$ and subtracting (4.6) we conclude that
$\left(\frac{2^{*}_{\eta_{1}}}{2}-1\right)$ $\int_{\mathbb{R}^{N}}|\nabla\tilde{u}_{n}|^{2}dx+\left(\frac{2^{*}_{\eta_{1}}}{q}-1\right)\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{q}}{|x|^{b}}dx$
$=\lambda\left(\frac{2^{*}_{\eta_{1}}}{p}-1\right)\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{p}}{|x|^{a}}dx+(\tilde{t}_{n})^{\ell_{2}-\ell}\mu\left(\frac{2^{*}_{\eta_{1}}}{r_{2}}-1\right)\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{r_{2}}}{|x|^{\eta_{2}}}dx+o(1).$ (4.7)
By the assumption on $\eta_{1}$ we have $2^{*}_{\eta_{1}}>q$ and $\ell_{1}>\ell$. Now, let us consider the conditions on $\mu$ and $(\eta_{2},r_{2})$.
[(i)] For $(\eta_{2},r_{2})$ satisfying i) we have $\ell_{2}\leq\ell$. Since the sequences $(\Phi_{\lambda}(\tilde{u}_{n}))$ and $(\|\tilde{u}_{n}\|_{L^{r_{2}}_{\eta_{2}}})$ are bounded and $\tilde{t}_{n}\to\infty$, it follows from (4.5) that $\tilde{u}_{n}\rightarrow 0$ in $L^{2^{*}_{\eta_{1}}}_{\eta_{1}}(\mathbb{R}^{N})$. This implies that the weak limit of $\tilde{u}_{n}$ in $E^{q}_{b}$ is $v\equiv 0$. Since $b<\eta_{2}$ and $1\leq r_{2}\in(2^{*}_{b,q,\eta_{2}},2^{*}_{\eta_{2}})$, the compactness of the embeddings of $E^{q}_{b}$ in $L^{r_{2}}_{\eta_{2}}(\mathbb{R}^{N})$ and $L^{p}_{a}(\mathbb{R}^{N})$ (see Lemma 2.2 and Corollary 2.3) give also
$\|\tilde{u}_{n}\|_{L^{r_{2}}_{\eta_{2}}}\rightarrow 0\quad\text{and}\quad\|\tilde{u}_{n}\|_{L^{p}_{a}}\rightarrow 0.$
Then, as $\ell_{2}\leq\ell$ and $2^{*}_{\eta_{1}}>q>2$, it follows from (4.1) that $\|\tilde{u}_{n}\|=o(1)$. Since $\tilde{u}_{n}\in\mathcal{M},$ this is a contradiction, and therefore the (PS) sequence $(u_{n})$ is bounded in $E^{q}_{b}$.
[(ii)] In this case, since $\mu>0$ (4.5) also yields $\tilde{u}_{n}\to 0$ in $L^{2^{*}_{\eta_{1}}}_{\eta_{1}}(\mathbb{R}^{N})$. Consequently $\tilde{u}_{n}\rightharpoonup 0$ in $E^{q}_{b}$ and so $\tilde{u}_{n}\to 0$ in $L^{p}_{a}(\mathbb{R}^{N})$. Once $2^{*}_{\eta_{1}}\leq r_{2}$ and $\mu>0$, by (4.1) we reach $\|\tilde{u}_{n}\|\leq o(1)$, which is an absurd.
[(iii)] Here we assume $\mu>0$, $\eta_{1}\leq\eta_{2}$ (equality only if $\eta_{1}=\eta_{2}>b$), and the conditions in $(\eta_{2},r_{2})$ ensures that $r_{2}<2^{*}_{\eta_{1}}$ and $\ell<\ell_{2}<\ell_{1}$. Then, from (4.5) we get
$\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx=O((\tilde{t}_{n})^{\ell-\ell_{1}}),\quad\text{as}\quad n\to\infty,$ (4.8)
which implies that $\tilde{u}_{n}\rightharpoonup 0$ in $E^{q}_{b}$. Let us denote
$\bar{\eta}:=\frac{(\eta_{1}-\eta_{2})(\delta q+b)-\delta\eta_{1}r_{2}+\delta\eta_{2}2^{*}_{\eta_{1}}}{\delta(2^{*}_{\eta_{1}}-r_{2})+\eta_{1}-\eta_{2}}\quad\text{and}\quad\bar{s}:=q+\frac{b-\bar{\eta}}{\delta}.$
Under the conditions in this item, we get $\bar{\eta}\in(b,2)$. Moreover, $\ell_{2}>\ell$ implies that $\bar{s}<r_{2}$. Let us show that $\bar{\eta}\in(b,2)$. Since $\ell_{1}>\ell_{2}$ we verify that
$\bar{\eta}\geq\eta_{2}\quad\Leftrightarrow\quad(\eta_{2}-\eta_{1})(\ell_{2}-\ell)\geq 0.$
Notice that, if $\eta_{1}=\eta_{2}>b$ we have $\bar{\eta}=\eta_{2}>b$. For $\eta_{2}\geq b$ and $\eta_{2}>\eta_{1}$, it holds $\bar{\eta}>\eta_{2}\geq b$. If $\eta_{1}<\eta_{2}<b$ we compare $\bar{\eta}$ directly with $b$, getting
$\bar{\eta}>b\quad\Leftrightarrow\quad(\eta_{2}-\eta_{1})(2^{*}_{\eta_{1}}-q)-(2^{*}_{\eta_{1}}-r_{2})(b-\eta_{1})>0\quad\Leftrightarrow\quad r_{2}>q+(2^{*}_{\eta_{1}}-q)\frac{b-\eta_{2}}{b-\eta_{1}},$
as we required. On the other hand, $\bar{\eta}<2$ iff $r_{2}<2^{*}_{\eta_{2}}$. Thus, for these $\bar{\eta}\in(b,2)$ and $\bar{s}=q+(b-\bar{\eta})/\delta$ we have $2<\bar{s}\in(2^{*}_{b,q,\bar{\eta}},2^{*}_{\bar{\eta}})$. Due to the compactness of the embeddings $E^{q}_{b}\hookrightarrow L^{\bar{s}}_{\bar{\eta}}(\mathbb{R}^{N})$ and $E^{q}_{b}\hookrightarrow L^{p}_{a}(\mathbb{R}^{N})$ we obtain
$\|\tilde{u}_{n}\|_{L^{\bar{s}}_{\bar{\eta}}}\rightarrow 0\quad\text{and}\quad\|\tilde{u}_{n}\|_{L^{p}_{a}}\rightarrow 0.$ (4.9)
Besides, by the choice of $\bar{\eta}$ and $\bar{s}$, we see that there exists $\theta\in(0,1)$ such that
$\frac{1}{r_{2}}=\frac{\theta}{2^{*}_{\eta_{1}}}+\frac{1-\theta}{\bar{s}}\quad\text{and}\quad\frac{\eta_{2}}{r_{2}}=\frac{\theta\eta_{1}}{2^{*}_{\eta_{1}}}+\frac{(1-\theta)\bar{\eta}}{\bar{s}}.$
Using the interpolation inequality in weighted Lebesgue spaces (see e.g. [7, Inequality (2.15)]) we get
$\left(\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{r_{2}}}{|x|^{\eta_{2}}}dx\right)^{\frac{1}{r_{2}}}\leq\left(\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx\right)^{\frac{\theta}{2^{*}_{\eta_{1}}}}\left(\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{\bar{s}}}{|x|^{\bar{\eta}}}dx\right)^{\frac{1-\theta}{\bar{s}}}.$
So, using (4.8) we obtain
$(\tilde{t}_{n})^{\ell_{2}-\ell}\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{r_{2}}}{|x|^{\eta_{2}}}dx\leq C(\tilde{t}_{n})^{\ell_{2}-\ell-(\ell_{1}-\ell)\frac{\theta r_{2}}{2^{*}_{\eta_{1}}}}\|\tilde{u}_{n}\|_{L^{\bar{s}}_{\bar{\eta}}}^{(1-\theta)r_{2}}=o(1),$
since $\ell_{2}-\ell=(\ell_{1}-\ell)(\theta r_{2}/2^{*}_{\eta_{1}})$. Adding this information to (4.9), from (4.1) we reach $\tilde{u}_{n}\to 0$ in $E^{q}_{b}$, which is an absurd because $\tilde{u}_{n}\in\mathcal{M}$.
Therefore, in each of the cases [(i)]-[(iii)] we have $(u_{n})$ a bounded sequence. ∎
Remark 4.1.
When $N\geq 3$, observe that $\max\{q,2^{*}_{b,q,\eta}\}=2^{*}_{b,q,\eta}$ whenever $\eta<b$, whereas it equals $q$ if $b\leq\eta<N$. Thus, conditions of the form $r>q$ for an admissible pair $(\eta,r)$ are automatically satisfied whenever $0\leq\eta<b$. Moreover, the condition $\eta<N-\frac{q(N-2)}{2}$ is equivalent to $q<2^{*}_{\eta}$. Since we assume $q<2^{*}_{b}$, this inequality also holds automatically for every $\eta<b$.
The next lemma provides the thresholds where one can recover the $(PS)_{c}$ conditions for the critical cases. As usual, this threshold is related to the constants of the Embeddings of $D^{1,2}$ into the Lebesgue spaces. Then, let us define, for $N\geq 3$ and each $0\leq\eta<2$,
$S_{\eta}=\inf_{u\in D^{1,2}(\mathbb{R}^{N})\setminus\{0\}}\frac{\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx}{\left(\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta}}}{|x|^{\eta}}\right)^{\frac{2}{2^{*}_{\eta}}}}.$ (4.10)
Lemma 4.3.
Assume ($\mathcal{H}$) and $N\geq 3$. Consider $f(x,u)$ as given in (1.9), with $\lambda\in\mathbb{R}$, $\mu\geq 0$, $\eta_{1}$ satisfying (1.10) and $(\eta_{2},r_{2})$ is a superscaled pair.
If $(r_{2},\eta_{2})$ satisfies (1.11) or (1.12), then the functional $\Phi$ defined in (1.6) satisfies $(PS)_{c}$ for all
$c<\frac{2-\eta_{1}}{2(N-\eta_{1})}S^{\frac{N-\eta_{1}}{2-\eta_{1}}}_{\eta_{1}}.$
If $(r_{2},\eta_{2})$ satisfies (1.13), define $\tilde{S}=\tilde{S}(\mu,\eta_{1},\eta_{2})>0$ such that $\mu S_{\eta_{2}}^{-\frac{2^{*}_{\eta_{2}}}{2}}\tilde{S}^{\frac{2-{\eta_{2}}}{N-2}}+S_{\eta_{1}}^{-\frac{2^{*}_{\eta_{1}}}{2}}\tilde{S}^{\frac{2-\eta_{1}}{N-2}}=1.$ Then, the functional $\Phi$ satisfies the $(PS)_{c}$ condition for all
$c<\frac{2-\max\{\eta_{1},\eta_{2}\}}{2(N-\max\{\eta_{1},\eta_{2}\})}\tilde{S}$.
Proof.
Let $c\in\mathbb{R}$ and let $(u_{n})$ be a $(PS)_{c}$ sequence for $\Phi$. We are under the hypotheses of Lemma 4.2. Therefore, $(u_{n})$ is bounded. Then, let $u\in E^{q}_{b}$ be such that $u_{n}\rightharpoonup u$ in $E^{q}_{b}$. By this weak convergence, we have
$\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx=\int_{\mathbb{R}^{N}}|\nabla u_{n}|^{2}dx-\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+o(1)$
and by Brezis-Lieb Lemma
$\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{q}}{|x|^{b}}dx=\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{q}}{|x|^{b}}dx-\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx+o(1)$
$\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx=\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx-\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx+o(1)$
with analogous identity for the pair $(\eta_{2},2^{*}_{\eta_{2}})$. For later use, notice that these relations imply that
$\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{q}}{|x|^{b}}dx\geq\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx+o(1)$ (4.11)
$\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{2^{*}_{\eta_{2}}}}{|x|^{\eta_{2}}}dx\geq\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{2}}}}{|x|^{\eta_{2}}}dx+o(1)$ (4.12)
Moreover, by the compactness of the embeddings
$E^{q}_{b}\hookrightarrow L^{p}_{a}(\mathbb{R}^{N}),\ E^{q}_{b}\hookrightarrow L^{r_{2}}_{\eta_{2}}(\mathbb{R}^{N})\quad(\text{if }r_{2}<2^{*}_{\eta_{2}}),$
we have $u_{n}\rightarrow u$ strongly in both spaces. Notice also that standard and well-known arguments show that $\Phi^{\prime}(u)v=0$ for all $v\in E^{q}_{b}$. Then, using all the convergences and Brezis-Lieb relations in the equation
$\Phi^{\prime}(u_{n})u_{n}-\Phi^{\prime}(u)u=o(1)$
gives, in the case $r_{2}<2^{*}_{\eta_{2}}$, the estimate
$\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx+\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{q}}{|x|^{b}}dx=\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx+o(1)$ (4.13)
or, in case $r_{2}=2^{*}_{\eta_{2}}$, the following
$\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx+\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{q}}{|x|^{b}}dx=\mu\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{2^{*}_{\eta_{2}}}}{|x|^{\eta_{2}}}dx+\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx+o(1).$ (4.14)
Proof of (i): Here we have $r_{2}<2^{*}_{\eta_{2}}$. We start from (4.13). Due to (4.10), we get
$\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx\leq\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx+\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{q}}{|x|^{b}}dx\leq S_{\eta_{1}}^{-\frac{2^{*}_{\eta_{1}}}{2}}\left(\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx\right)^{\frac{2^{*}_{\eta_{1}}}{2}}\!\!\!+o(1).$
Assume, by contradiction, that $(u_{n})$ does not converge to $u$ in $E_{b}^{q}$, not even along a subsequence. Then, by (4.13) and the last inequality above, the same lack of convergence holds in $D^{1,2}(\mathbb{R}^{N})$. So we can divide the last inequality by $\|\nabla(u_{n}-u)\|_{L^{2}}^{2}$ and get
$\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx\geq S_{\eta_{1}}^{\frac{N-\eta_{1}}{2-\eta_{1}}}+o(1).$
Then,
$\int_{\mathbb{R}^{N}}|\nabla u_{n}|^{2}dx-\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx=\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx+o(1)\geq S_{\eta_{1}}^{\frac{N-\eta_{1}}{2-\eta_{1}}}+o(1).$ (4.15)
Now, divide the equation $\Phi^{\prime}(u_{n})u_{n}=o(1)$ by $2^{*}_{\eta_{1}}$ and subtract if from $\Phi(u_{n})=c+o(1)$. This gives
$c+o(1)\,\,=\,\,$ $\frac{2-\eta_{1}}{2(N-\eta_{1})}\int_{\mathbb{R}^{N}}|\nabla u_{n}|^{2}dx+\left(\frac{1}{q}-\frac{1}{2^{*}_{\eta_{1}}}\right)\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{q}}{|x|^{b}}dx$
$-\lambda\left(\frac{1}{p}-\frac{1}{2^{*}_{\eta_{1}}}\right)\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{p}}{|x|^{a}}dx-\mu\left(\frac{1}{r_{2}}-\frac{1}{2^{*}_{\eta_{1}}}\right)\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{r_{2}}}{|x|^{\eta_{2}}}dx.$
Now, recall (4.11) and (4.15) and let $n\rightarrow\infty$ in this last equality to reach
$c\,\,\geq\,\,$ $\frac{2-\eta_{1}}{2(N-\eta_{1})}S_{\eta_{1}}^{\frac{N-\eta_{1}}{2-\eta_{1}}}+\frac{2-\eta_{1}}{2(N-\eta_{1})}\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\left(\frac{1}{q}-\frac{1}{2^{*}_{\eta_{1}}}\right)\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx$
$-\lambda\left(\frac{1}{p}-\frac{1}{2^{*}_{\eta_{1}}}\right)\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx-\mu\left(\frac{1}{r_{2}}-\frac{1}{2^{*}_{\eta_{1}}}\right)\int_{\mathbb{R}^{N}}\frac{|u|^{r_{2}}}{|x|^{\eta_{2}}}dx,$ (4.16)
where we have used that $u_{n}\rightarrow u$ both in $L^{p}_{a}(\mathbb{R}^{N})$ and $L^{r_{2}}_{\eta_{2}}(\mathbb{R}^{N})$. We need now to invoke the Pohozaev Identity given in (2.7) (here is where we need to assume $\eta_{1}<4/N$). Multiply the Pohozaev Identity by $\alpha$ and $\Phi^{\prime}(u)u=0$ by $\beta$, where $\alpha$ and $\beta$ will be chosen later. Then, add the results in (4.16). We get
$c\,\,\geq\,\,$ $\frac{2-\eta_{1}}{2(N-\eta_{1})}S_{\eta_{1}}^{\frac{N-\eta_{1}}{2-\eta_{1}}}\!+\left[\frac{2-\eta_{1}}{2(N-\eta_{1})}+\frac{N-2}{2}\alpha+\beta\right]\!\int_{\mathbb{R}^{N}}\!\!|\nabla u|^{2}dx$
$+\left[\left(\frac{1}{q}-\frac{1}{2^{*}_{\eta_{1}}}\right)\!+\frac{N-b}{q}\alpha+\beta\right]\!\int_{\mathbb{R}^{N}}\!\!\frac{|u|^{q}}{|x|^{b}}dx-\lambda\left[\left(\frac{1}{p}-\frac{1}{2^{*}_{\eta_{1}}}\right)+\frac{N-a}{p}\alpha+\beta\right]\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx$
$-\mu\left[\left(\frac{1}{r_{2}}-\frac{1}{2^{*}_{\eta_{1}}}\right)+\frac{N-\eta_{2}}{r_{2}}\alpha+\beta\right]\int_{\mathbb{R}^{N}}\frac{|u|^{r_{2}}}{|x|^{\eta_{2}}}dx-\left[\frac{N-2}{2}\alpha+\beta\right]\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}.$ (4.17)
Notice that we used that $(N-\eta_{1})/2^{*}_{\eta_{1}}=(N-2)/2$. Now choose
$\alpha=\frac{q-2}{2(N-b)-q(N-2)}\qquad\text{and}\qquad\beta=-\frac{2-\eta_{1}}{2(N-\eta_{1})}-\frac{N-2}{2}\,\alpha.$
With this choice, the first three expressions in $\cdots$] vanish, while the last one becomes strictly negative. It remains to verify that
$\left(\frac{1}{r_{2}}-\frac{1}{2^{*}_{\eta_{1}}}\right)+\frac{N-\eta_{2}}{r_{2}}\,\alpha+\beta<0,$
which is equivalent to
$r_{2}>2\,\frac{1+\alpha(N-\eta_{2})}{1+\alpha(N-2)}.$
However, this is a requirement for the pair $(\eta_{2},r_{2})$ to be called superscaled, since
$2\frac{1+\alpha(N-\eta_{2})}{1+\alpha(N-2)}=q+\frac{b-\eta_{2}}{\delta}.$
Finally, we go back to (4.17) and see that
$c\geq\frac{2-\eta_{1}}{2(N-\eta_{1})}S_{\eta_{1}}^{\frac{N-\eta_{1}}{2-\eta_{1}}}.$
This completes the proof for the item (i).
Proof of (ii): Now, let us consider $r_{2}=2^{*}_{\eta_{2}}$ and $\eta_{2}$ satisfying (1.10). From (4.14), using (4.10), we get
$\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx\leq\mu S_{\eta_{2}}^{-\frac{2^{*}_{\eta_{2}}}{2}}\left(\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx\right)^{\frac{2^{*}_{\eta_{2}}}{2}}+S_{\eta_{1}}^{-\frac{2^{*}_{\eta_{1}}}{2}}\left(\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx\right)^{\frac{2^{*}_{\eta_{1}}}{2}}\!\!\!+o(1).$
Here, analogously to the previous case, we suppose that $(u_{n})$ does not have any subsequence converging to $u$ in $E^{q}_{b}$, which, due to (4.14), implies that $\|\nabla(u_{n}-u)\|^{2}_{L^{2}}$ stays away from 0 for large values of $n$. So, we can divide the last inequality by $\|\nabla(u_{n}-u)\|^{2}_{L^{2}}$ and get
$\mu S_{\eta_{2}}^{-\frac{2^{*}_{\eta_{2}}}{2}}\left(\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx\right)^{\frac{2-\eta_{2}}{N-\eta_{2}}}+S_{\eta_{1}}^{-\frac{2^{*}_{\eta_{1}}}{2}}\left(\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx\right)^{\frac{2-\eta_{1}}{N-\eta_{1}}}\geq 1+o(1).$
Thus, the definition of $\tilde{S}$ in the assumptions of this lemma gives us
$\int_{\mathbb{R}^{N}}|\nabla u_{n}|^{2}dx-\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx=\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx+o(1)\geq\tilde{S}+o(1).$
Now, without loss of generality, suppose $\eta_{1}\geq\eta_{2}$. The other case can be handled by just interchanging their roles in the following arguments. We divide the identity $\Phi^{\prime}(u_{n})u_{n}=o(1)$ by $2^{*}_{\eta_{1}}$ and subtract it from $\Phi(u_{n})=c+o(1)$. This yields
$c+o(1)\,\,=\,\,$ $\frac{2-\eta_{1}}{2(N-\eta_{1})}\int_{\mathbb{R}^{N}}|\nabla u_{n}|^{2}dx+\left(\frac{1}{q}-\frac{1}{2^{*}_{\eta_{1}}}\right)\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{q}}{|x|^{b}}dx$
$-\lambda\left(\frac{1}{p}-\frac{1}{2^{*}_{\eta_{1}}}\right)\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{p}}{|x|^{a}}dx+\left(\frac{1}{2^{*}_{\eta_{1}}}-\frac{1}{2^{*}_{\eta_{2}}}\right)\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{2^{*}_{\eta_{2}}}}{|x|^{\eta_{2}}}dx.$
By assumption, we have $q<2^{*}_{\eta_{1}}\leq 2^{*}_{\eta_{2}}$ . So, we use (4.11) and (4.12) and letting $n\rightarrow\infty$ in the last inequality, we have
$c\,\,\geq\,\,$ $\frac{2-\eta_{1}}{2(N-\eta_{1})}\tilde{S}+\frac{2-\eta_{1}}{2(N-\eta_{1})}\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\left(\frac{1}{q}-\frac{1}{2^{*}_{\eta_{1}}}\right)\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx$
$-\lambda\left(\frac{1}{p}-\frac{1}{2^{*}_{\eta_{1}}}\right)\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx+\left(\frac{1}{2^{*}_{\eta_{1}}}-\frac{1}{2^{*}_{\eta_{2}}}\right)\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{2}}}}{|x|^{\eta_{2}}}dx.$
We now proceed exactly as in the subcritical case $r_{2}<2^{*}_{\eta_{2}}$. This leads to the conclusion that
$c\,\,\geq\,\,\frac{2-\eta_{1}}{2(N-\eta_{1})}\tilde{S}$,
and, of course, if $\eta_{2}>\eta_{1}$ we would have $\eta_{2}$ instead of $\eta_{1}$ in this last estimate. This finishes this proof. ∎
Remark 4.2.
The $(PS)_{c}$ condition also holds for levels below the threshold in Lemma 4.3 when $(r_{2},\eta_{2})$ is subscaled, $b<\eta_{2}<2$ and $\mu\leq 0$. Indeed, notice that $(PS)_{c}$ sequences are bounded due to Lemma 4.2. To prove that there exists convergent subsequences, recall that in Lemma 4.3 (the case with only one critical pair) the argument required
$\mu\left[\left(\frac{1}{r_{2}}-\frac{1}{2^{*}_{\eta_{1}}}\right)+\frac{N-\eta_{2}}{r_{2}}\,\alpha+\beta\right]\leq 0,$
for the specific choices of $\alpha$ and $\beta$ made there. That inequality was valid because $\mu\geq 0$ and $(\eta_{2},r_{2})$ was superscaled. The same inequality remains valid when $\mu\leq 0$ and $(\eta_{2},r_{2})$ is subscaled, and thus the $(PS)_{c}$ condition holds in this setting as well.
4.2 Existence of solutions for subcritical superscaled problems
Here we prove that (1.4) has a nontrivial solution for subcritical superscaled nonlinearities. We present distinct proofs for Theorem 1.1, concerning $\lambda\in\sigma(\mathcal{A},\mathcal{B})$ or not. If $\lambda$ is an eigenvalue for (1.1), we apply a recent abstract result based on a notion of local linking (see [23, Theorem 2.30 and Corollary 2.31]).
Proofof Theorem 1.1. Lemma 4.1 ensures that $\Phi$ satisfies the (PS) condition. For $G(x,t)=|x|^{-\eta}H(t)$, let us denote $\tilde{F},\tilde{G}:E^{q}_{b}\to\mathbb{R}$ by
$\tilde{F}(u)=\int_{\mathbb{R}^{N}}F(x,u)dx=\frac{\lambda}{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx+\int_{\mathbb{R}^{N}}\frac{H(u)}{|x|^{\eta}}dx\quad\text{and}\quad\tilde{G}(u)=\int_{\mathbb{R}^{N}}\frac{H(u)}{|x|^{\eta}}dx,$
and $\tilde{g}=\tilde{G}^{\prime}$. For $\ell_{i}=\ell(\eta_{i},r_{i})=r_{i}\delta+\eta-N$, $i=1,2$, due to (1.8), we verify that
$|\tilde{G}(u_{t})|$ $\leq\frac{c_{1}t^{\,\ell_{1}}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{1}}}{|x|^{\eta}}dx+\frac{c_{2}t^{\,\ell_{2}}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{2}}}{|x|^{\eta}}dx,$
$|\tilde{g}(u_{t})v_{t}|$ $\leq c_{1}t^{\,\ell_{1}}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{1}-1}|v|}{|x|^{\eta}}dx+{c_{2}t^{\,\ell_{2}}}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{2}-1}|v|}{|x|^{\eta}}dx\leq c_{1}t^{\,\ell_{1}}\|u\|_{L^{r_{1}}_{\eta}}^{r_{1}-1}\|v\|_{L^{r_{1}}_{\eta}}+c_{2}t^{\,\ell_{2}}\|u\|_{L^{r_{2}}_{\eta}}^{r_{2}-1}\|v\|_{L^{r_{2}}_{\eta}}$
and
$\tilde{F}(u_{t})\geq\tilde{G}(u_{t})\geq\frac{ct^{\,\ell(\eta,r)}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx.$
Using the embeddings of $E^{q}_{b}$ in $L^{r_{i}}_{\eta}(\mathbb{R}^{N})$ and that $\ell<{\ell(\eta,r)},\ell_{i}$ (since we considered superscaled pairs), $i=1,2$, we obtain $\tilde{g}(u_{t})v_{t}=o(t^{\,\ell})\|v\|$ and $\tilde{G}(u_{t})=o(t^{\,\ell})$ as $t\to 0^{+}$}, uniformly in $u$ on bounded subsets, and
$\lim_{t\to\infty}\frac{\tilde{F}(u_{t})}{t^{\,\ell}}=\infty,$
uniformly in $u$ on compact subsets of $\mathcal{M}.$ Then, if $\lambda\notin\sigma(\mathcal{A},\mathcal{B})$, we can apply [23, Theorem 2.27] to guarantee the existence of a nontrivial solution for (1.4) at a positive level. In particular, since $\sigma(\mathcal{A},\mathcal{B})\subset[\lambda_{1},\infty)$, it holds for any $\lambda<\lambda_{1}$.
In general, for $\lambda\geq\lambda_{1}$ (possibly an eigenvalue) we will apply [23, Corollary 2.31] to prove this result. Consider $k\in\mathbb{N}$ such that $\lambda_{k}\leq\lambda<\lambda_{k+1}$. Our first step is to show that $\Phi$ has a scaled local linking near the origin in dimension $k$, which means that there exist two nonempty, closed, symmetric and disjoint subsets $A_{0},B_{0}\subset\mathcal{M}$ such that
$i(A_{0})=i(\mathcal{M}\backslash B_{0})=k$
and $\rho>0$ satisfying
$\begin{cases}\Phi(u_{t})\leq 0,&\forall u\in A_{0}\,\,\mbox{and}\,\,0\leq t\leq\rho,\\ \Phi(u_{t})>0,&\forall u\in B_{0}\,\,\mbox{and}\,\,0<t\leq\rho.\end{cases}$ (4.18)
Consider $A_{0}=\tilde{\Psi}^{\lambda_{k}}$ and $B_{0}=\tilde{\Psi}_{\lambda_{k+1}}$. Since $\lambda_{k}<\lambda_{k+1}$, by Theorem 3.3(iii) we have
$i(\tilde{\Psi}^{\lambda_{k}})=i(\mathcal{M}\backslash\tilde{\Psi}_{\lambda_{k+1}})=k.$
Moreover, for $u\in B_{0}$ we have $\tilde{\Psi}(u)\geq\lambda_{k+1}$ so that $J(u)\leq 1/\lambda_{k+1}$ and then
$\Phi(u_{t})=I(u_{t})-\lambda J(u_{t})-\tilde{G}(u_{t})\geq t^{\,\ell}\left(1-\frac{\lambda}{\lambda_{k+1}}+o(1)\right)\quad\text{as}\,\,t\to 0^{+},$
uniformly on $u\in\mathcal{M}$. Thus, since $\lambda<\lambda_{k+1}$, for small $\rho>0$ we see that $\Phi$ satisfies (4.18) on $B_{0}$. On the other hand, for $u\in A_{0}$ we have $\tilde{\Psi}(u)\leq\lambda_{k}\leq\lambda$. As $\tilde{G}\geq 0$, for all $u\in A_{0}$ and $t\in[0,\rho]$ we get
$\Phi(u_{t})=I(u_{t})-\lambda J(u_{t})-\tilde{G}(u_{t})\leq t^{\,\ell}\left(1-\frac{\lambda}{\lambda_{k}}\right)\leq 0.$
Thus (4.18) holds and $\Phi$ has a scaled local linking near the origin in dimension $k$.
Next we need to show that $\Phi^{a}$ is contractible for some $\alpha<0$. Consider
$\alpha<\inf_{u\in\mathcal{M},\,0\leq t\leq 1}\Phi(u_{t}).$
Denoting $\varphi_{u}(t)=\Phi(u_{t})$ we have
$\Phi^{\alpha}=\{u\in E^{q}_{b}:\Phi(u)\leq\alpha\}=\{u_{t}:u\in\mathcal{M},\,t>1\,\,\mbox{and}\,\,\varphi_{u}(t)\leq\alpha\}.$
Hence, using again (1.8), for each $u\in\mathcal{M}$ we get
$\varphi_{u}(t)$ $=$ $I(u_{t})-\lambda J(u_{t})-\tilde{G}(u_{t})=t^{\,\ell}\left(1-{\lambda}J(u)\right)-\int_{\mathbb{R}^{N}}\frac{H(u_{t})}{|x|^{\eta}}dx$
$\leq$ $t^{\,\ell}\left(1-\frac{\lambda}{\tilde{\Psi}(u)}\right)-\frac{ct^{{\ell(\eta,r)}}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx\,\,\to\,\,-\infty\quad\text{as}\,\,t\to\infty,$
since ${\ell(\eta,r)}>\ell>0$. Notice that
$\int_{\mathbb{R}^{N}}\frac{H(u_{t})}{|x|^{\eta}}dx=\int_{\mathbb{R}^{N}}\frac{H(t^{\delta}u(tx))}{|x|^{\eta}}dx=t^{\eta-N}\int_{\mathbb{R}^{N}}\frac{H(t^{\delta}u(x))}{|x|^{\eta}}dx$
and so $\varphi_{u}$ is a $C^{1}$ function satisfying
$t\varphi^{\prime}_{u}(t)$ $=\ell t^{\,\ell}\left(1-\frac{\lambda}{\tilde{\Psi}(u)}\right)-(\eta-N)t^{\eta-N}\int_{\mathbb{R}^{N}}\frac{H(t^{\delta}u)}{|x|^{\eta}}dx-\delta t^{\eta-N}\int_{\mathbb{R}^{N}}\frac{h(t^{\delta}u)t^{\delta}u}{|x|^{\eta}}dx$
$\leq\ell t^{\,\ell}\left(1-\frac{\lambda}{\tilde{\Psi}(u)}\right)-\ell(\eta,r)t^{\eta-N}\int_{\mathbb{R}^{N}}\frac{H(t^{\delta}u)}{|x|^{\eta}}dx\,\,\leq\,\,\ell\varphi_{u}(t),\quad\mbox{for all}\,\,t>0.$
So, $\varphi_{u}(t)<0$ yields $\varphi^{\prime}_{u}(t)<0$. Then, the implicit function theorem provides a $C^{1}$-map $\psi:\mathcal{M}\to(1,\infty)$ such that
$\varphi_{u}(t)>\alpha\,\,\mbox{for}\,\,0\leq t<\psi(u),\quad\varphi_{u}(\psi(u))=\alpha\quad\mbox{and}\quad\varphi_{u}(t)<\alpha\,\,\forall t>\psi(u),$
so that
$\Phi^{\alpha}=\{u_{t}:u\in\mathcal{M},\,t\geq\psi(u)\}.$
Finally, for $t_{u}$ and $\tilde{u}=\pi(u)=u_{t_{u}}$ as in (3.7), we consider $T:(E\backslash\{0\})\times[0,1]\to E\backslash\{0\}$ defined as
$T(u,\theta)=\begin{cases}\tilde{u}_{\theta\psi(\tilde{u})+(1-\theta)t^{-1}_{u}}&\mbox{if}\,\,u\notin\Phi^{\alpha}\\ u&\mbox{if}\,\,u\in\Phi^{\alpha},\end{cases}$
[…]
Next we are going to prove our multiplicity result for subcritical problems.
Proofof Theorem 1.2. To prove this result apply [25, Propositions 3.42 and 3.44]. From Lemma 4.1 we know that $\Phi$ satisfies the $(PS)_{c}$ condition for all $c\in\mathbb{R}$, in particular for $c\in(0,\infty)$. Since we are assuming that $h$ is an odd function we have $\Phi$ an even functional. We set
$B_{0}:=\tilde{\Psi}_{\lambda_{k_{0}}}\quad\mbox{for}\quad k_{0}:=\min\{k\in\mathbb{N}:\lambda<\lambda_{k}\}.$
If $k_{0}=1$, by Theorem 3.3(i) we have
$B_{0}=\{u\in\mathcal{M}:\tilde{\Psi}(u)\geq\lambda_{1}\}=\mathcal{M}.$
Hence $i(\mathcal{M}\backslash B_{0})=i(\emptyset)=0=k_{0}-1$. If $k_{0}\geq 2$, it holds that $\lambda_{k_{0}-1}\leq\lambda<\lambda_{k_{0}}$, hence Theorem 3.3(iii) yields
$i(\mathcal{M}\backslash B_{0})=i(\mathcal{M}\backslash\tilde{\Psi}_{\lambda_{k_{0}}})=k_{0}-1.$
For $k\geq k_{0}$ in $\mathbb{N}$ we consider $W_{k}$ a subspace of $E^{q}_{b}$ with dimension $k$ and denote $S^{k-1}$ the unit sphere of $W_{k}$, which is a compact set in $E^{q}_{b}$ and satisfies $i(S^{k-1})=k$ (see [12]). For $R,\rho>0$ to be chosen later, we set
$B=\{u_{\rho}:u\in B_{0}\},\quad A=\{Ru:u\in S^{k-1}\}\quad\mbox{and}\quad X=\{tu:u\in A,\,0\leq t\leq 1\}.$
We will verify that we may choose $R>\rho>0$ such that $A\subset E^{q}_{b}\setminus I^{-1}[0,\rho]$ and
$\sup_{v\in A}\Phi(v)\leq 0<\inf_{v\in B}\Phi(v)\quad\mbox{and}\quad\sup_{v\in X}\Phi(v)<\infty.$ (4.19)
From (1.8) we have $|H(s)|\leq c_{1}|s|^{r_{1}}+c_{2}|s|^{r_{2}}$ for all $s\in\mathbb{R}$, for superscaled pairs $(\eta,r_{i})$. Then,
$\tilde{G}(u_{t})=\int_{\mathbb{R}^{N}}\frac{H(u_{t})}{|x|^{\eta}}dx=o(t^{\,\ell}),\quad\text{as}\,\,\,t\to 0^{+},$
uniformly in $u\in B_{0}$. See the proof of Theorem 1.1 for details. Since $(J(u))^{-1}=\tilde{\Psi}(u)\geq\lambda_{k_{0}}$ in $B_{0}$ we get
$\Phi(u_{t})$ $={t^{\,\ell}}I(u)-\lambda t^{\,\ell}J(u)-\tilde{G}(u_{t})\geq t^{\,\ell}\left(1-\frac{\lambda^{+}}{\lambda_{k_{0}}}+o(1)\right),\quad\text{as}\,\,\,t\to 0^{+}.$
Once $\lambda^{+}=\max\{\lambda,0\}<\lambda_{k_{0}}$, for $t=\rho>0$ sufficiently small it holds
$\inf_{v\in B}\Phi(v)=\inf_{u\in B_{0}}\Phi(u_{\rho})>0.$ (4.20)
On the other hand, due to (1.8) we have $rH(s)\geq c|s|^{r}$ for $s\in\mathbb{R}$. So
$\Phi(tu)$ $\leq\frac{t^{2}}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\frac{t^{q}}{q}\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx-\frac{\lambda t^{p}}{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx-\frac{ct^{r}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx.$ (4.21)
The compactness of $S^{k-1}$, and its boundedness in $L^{p}_{a}(\mathbb{R}^{N})$ and in $L^{r}_{\eta}(\mathbb{R}^{N})$, ensures that
$\frac{c}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx\geq c_{0}\quad\text{and}\quad\frac{|\lambda|}{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx\leq\tilde{c}_{1},\quad\forall u\in S^{k-1},$
for some constants $c_{0},\tilde{c}_{1}>0$. Recalling that $r>q>p>2$, from (4.21) we obtain
$\Phi(tu)\leq\frac{t^{2}}{2}+\frac{t^{q}}{q}+\tilde{c}_{1}t^{p}-c_{0}t^{r}\leq 0,\quad\forall u\in S^{k-1},$
whether $t=R$ is large enough, yielding also $A=RS^{k-1}\subset E^{q}_{b}\setminus I^{-1}[0,\rho]$. Hence, we get
$\sup_{v\in A}\Phi(v)=\sup_{u\in S^{k-1}}\Phi(Ru)\leq 0.$ (4.22)
Moreover,
$\sup_{v\in X}\Phi(v)=\sup_{t\in[0,1],\,u\in A}\Phi(tu)\leq\max_{t\geq 0}\left(\frac{t^{2}}{2}+\frac{t^{q}}{q}+\tilde{c}_{1}t^{p}-c_{0}t^{r}\right)<\infty.$
Joining this estimate with (4.20) and (4.22) we get (4.19). Therefore, using [25, Propositions 3.42 and 3.44] we obtain a sequence of positive critical values $(c^{*}_{k})$, $k\geq k_{0}$}, such that $c^{*}_{k}\nearrow\infty$. $\Box$
4.3 Existence and multiplicity for critical problems
In the proof of our results, we need some estimates on the norm of $u$ in the weighted Lebesgue spaces $L^{r}_{\eta}$, and we present it in the next lemmas.
Lemma 4.4.
[…]
$\left(\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}\,dx\right)^{\frac{1}{p}}\leq\left(\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}\,dx\right)^{\frac{\theta}{r}}\left(\int_{\mathbb{R}^{N}}\frac{|u|^{\tilde{r}}}{|x|^{\tilde{\eta}}}\,dx\right)^{\frac{1-\theta}{\tilde{r}}}$. (4.23)
[…]
Proof.
We use interpolation inequality with weights. If $\eta=a$ and $r\neq p$ the conclusion is immediate: choose $\tilde{\eta}=a$, $\max\{2^{*}_{b,q,a},1\}<\tilde{r}<2^{*}_{a}$ and $0<\theta<1$ such that
$\frac{1}{p}=\frac{\theta}{r}+\frac{1-\theta}{\tilde{r}}$
[…]
$\frac{1}{p}=\frac{\theta}{r}+\frac{1-\theta}{\tilde{r}},\qquad\frac{a}{p}=\frac{\theta\eta}{r}+\frac{(1-\theta)\tilde{\eta}}{\tilde{r}}$. (4.24)
These identities give
$\theta(\tilde{\eta})=\frac{a-\tilde{\eta}}{p}\,\frac{r}{\eta-\tilde{\eta}}\qquad\text{and}\qquad\frac{1}{\tilde{r}}=\frac{1}{p}+\frac{\theta(\tilde{\eta})}{1-\theta(\tilde{\eta})}\biggl(\frac{1}{p}-\frac{1}{r}\biggr).$ (4.25)
We distinguish two situations:
Case (i): $\eta>a$. Pick $\tilde{\eta}<a$, and so, $\eta-\tilde{\eta}>\eta-a>0$. Then letting $\tilde{\eta}\nearrow a$ we see from (4.25) that
$\theta(\tilde{\eta})\searrow 0,\qquad 2^{*}_{b,q,\tilde{\eta}}\searrow 2^{*}_{b,q,a},\qquad 2^{*}_{\tilde{\eta}}\searrow 2^{*}_{a}\qquad\tilde{r}\to p.$
Since $2^{*}_{b,q,a}<p<2^{*}_{a}$, we may choose $\tilde{\eta}$ sufficiently close to $a\in(b,2)$ so that
$b<\tilde{\eta}<2,\qquad 0<\theta(\tilde{\eta})<1,\qquad\max\{1,2^{*}_{b,q,\tilde{\eta}}\}<\tilde{r}<2^{*}_{\tilde{\eta}}$.
Case (ii): $\eta<a$. The argument is completely analogous: now select $\tilde{\eta}>a$, sufficiently close to $a$ so $b<\tilde{\eta}<2$, $0<\theta(\tilde{\eta})<1$ and $2^{*}_{b,q,\tilde{\eta}}<\tilde{r}<2^{*}_{\tilde{\eta}}$.
[…]
$\delta\left(\frac{p+(a-\tilde{\eta})/\delta}{\tilde{r}}-1\right)=\frac{(a-\tilde{\eta})}{p(\eta-a)+(a-\tilde{\eta})(p-r)}\,(\ell(\eta,r)-\ell\,)=:K(\tilde{\eta})\big(\ell(\eta,r)-\ell\,\big).$
[…]
Finally, using the interpolation inequality with distinct weights (see e.g. [7, Inequality (2.15)]), the relation (4.24) yields (4.23). ∎
Using the interpolation inequality from Lemma 4.4 we get some uniform estimate in $L^{r}_{\eta}(\mathbb{R}^{N})$ for functions in the set levels $\tilde{\Psi}^{\lambda}$.
Lemma 4.5.
Suppose $(\eta,r)$ is an admissible pair. Then, for each $\lambda\geq\lambda_{1}$ there exists a positive constant $c$, which depends on $\lambda,\eta,r$ and other fixed parameters, such that
$\frac{1}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx\geq c,\quad\forall u\in\tilde{\Psi}^{\lambda}.$ (4.26)
Proof.
[…]
$\frac{1}{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx\leq\frac{1}{p}\left(\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx\right)^{\frac{p\theta}{r}}\left(\int_{\mathbb{R}^{N}}\frac{|u|^{\tilde{r}}}{|x|^{\tilde{\eta}}}dx\right)^{\frac{p(1-\theta)}{\tilde{r}}}\leq C\left(\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx\right)^{\frac{\theta}{r}},$
for some positive constant $C=C(N,b,q,p,\tilde{\eta},\tilde{r},\theta)$. Since (4.26) holds for $(\eta,r)=(a,p)$, we conclude that it occurs for any admissible pair $(\eta,r)$. ∎
For reader’s convenience we recall the definition of pseudo-index. For any $\rho>0$, we define
$\mathcal{M}_{\rho}=\{u\in E^{q}_{b}:I(u)=\rho^{\ell}\}=\{u_{\rho}:u\in\mathcal{M}\}.$
Let $\Gamma$ be the group of odd homeomorphisms of $E^{q}_{b}$ that act as the identity outside the set $\Phi^{-1}(0,c^{*})$, for some $c^{*}>0$ such that $\Phi$ satisfies the $(\mathrm{PS})_{c}$ condition for all $c\in(0,c^{*})$. Let $\mathcal{A}^{*}$ be the family of symmetric subsets of $E^{q}_{b}$. For any $M\in\mathcal{A}^{*}$, the pseudo-index $i^{*}(M)$ with respect to $i$, $\mathcal{M}_{\rho}$, and $\Gamma$ (see Benci [6]) is defined
$i^{*}(M)=\min_{\gamma\in\Gamma}i(\gamma(M)\cap\mathcal{M}_{\rho}).$
The proof of Theorem 1.3 and Theorem 1.5 will be based on the following
Proposition 4.6 ([23], Theorem 2.33).
Let $A_{0}$ and $B_{0}$ be symmetric subsets of $\mathcal{M}$ such that $A_{0}$ is compact, $B_{0}$ is closed, and
$i(A_{0})\geq k+m-1,\qquad i(\mathcal{M}\setminus B_{0})\leq k-1$
for some $k,m\geq 1$. Let $R>\rho>0$ and let
$X=\{u_{t}:u\in A_{0},\,0\leq t\leq R\},\quad A=\{u_{R}:u\in A_{0}\},\quad B=\{u_{\rho}:u\in B_{0}\}.$
Assume that
$\sup_{u\in A}\Phi(u)\leq 0<\inf_{u\in B}\Phi(u),\qquad\sup_{u\in X}\Phi(u)<c^{*}.$
For $j=k,\ldots,k+m-1$, let
$\mathcal{A}^{*}_{j}=\{M\in\mathcal{A}^{*}:M\text{ is compact and }i^{*}(M)\geq j\}$
and set
$c^{*}_{j}:=\inf_{M\in\mathcal{A}^{*}_{j}}\max_{u\in M}\Phi(u).$
Then $0<c^{*}_{k}\leq\cdots\leq c^{*}_{k+m-1}<c^{*}$, each $c^{*}_{j}$ is a critical value of $\Phi$, and $\Phi$ has $m$ distinct pairs of associated critical points.
The next abstract result is a direct corollary of the preceding one and will be used in the proofs of Theorem 1.10.
Corollary 4.7 ([23], Corollary 2.34).
Let $A_{0}$ be a compact symmetric subset of $\mathcal{M}$ with $i(A_{0})=m\geq 1$, let $R>\rho>0$, and let
$A=\{u_{R}:u\in A_{0}\},\qquad X=\{u_{t}:u\in A_{0},\,0\leq t\leq R\}.$
Assume that
$\sup_{u\in A}\Phi(u)\leq 0<\inf_{u\in\mathcal{M}_{\rho}}\Phi(u),\qquad\sup_{u\in X}\Phi(u)<c^{*}.$
Then $\Phi$ has $m$ distinct pairs of critical points at levels in $(0,c^{*})$.
Now we prove the result of multiplicity of solution for $\lambda$ close to some eigenvalue $\lambda_{k}.$
Proofof Theorem 1.3. To prove this theorem we will apply Proposition 4.6. Let us then consider $\lambda\in(0,\lambda_{k})$. By Lemma 4.3 we know that $\Phi$ satisfies the $(PS)_{c}$ for all $c<c^{*}$, for a positive constant $c^{*}$ depending on $N$ and $\eta_{1}$ (and $\mu,\eta_{2}$ if $r_{2}=2^{*}_{\eta_{2}}$). Fix $\varepsilon\in(0,\lambda_{k+m}-\lambda_{k+m-1})$. Then, by Theorem 3.3(iii) we have
$i(\mathcal{M}\backslash\tilde{\Psi}_{\lambda_{k+m-1}+\varepsilon})=k+m-1.$
Since $\mathcal{M}\backslash\tilde{\Psi}_{\lambda_{k+m-1}+\varepsilon}$ is an open symmetric subset of $\mathcal{M}$, there is a compact symmetric $C_{\varepsilon}\subset\mathcal{M}\backslash\tilde{\Psi}_{\lambda_{k+m-1}+\varepsilon}$ with $i(C_{\varepsilon})=k+m-1$ (see the proof of Proposition 3.1 in Degiovanni and Lancelotti [11]). We set $A_{0}=C_{\varepsilon}$ and $B_{0}=\tilde{\Psi}_{\lambda_{k}}$. The properties of $C_{\varepsilon}$ yield
$i(A_{0})=k+m-1.$
As far as $B_{0},$ we distinguish the cases $\lambda_{k}=\lambda_{1}$ and $\lambda_{k}>\lambda_{1}$. If $\lambda_{1}=\cdots=\lambda_{k}$, we have
$B_{0}=\{u\in\mathcal{M}:\tilde{\Psi}(u)\geq\lambda_{1}\}=\mathcal{M},$
by Theorem 3.3(i). Then, $i(\mathcal{M}\backslash B_{0})=i(\emptyset)=0\leq k-1$. In the other case, it holds $\lambda_{l-1}<\lambda_{l}=\cdots=\lambda_{k}$ for some $2\leq l\leq k$. Then, Theorem 3.3(iii) implies that
$i(\mathcal{M}\backslash B_{0})=i(\mathcal{M}\backslash\tilde{\Psi}_{\lambda_{l}})=l-1\leq k-1.$
Now, for some $R>\rho>0$, we set
$X=\{u_{t}:u\in A_{0},\,0\leq t\leq R\},\quad A=\{u_{R}:u\in A_{0}\}\quad\mbox{and}\quad B=\{u_{\rho}:u\in B_{0}\},$
and show that it is possible to choose suitable $R,\rho$ and $\delta_{k}>0$ in such a way that, for all $\lambda\in(\lambda_{k}-\delta_{k},\lambda_{k})$, it holds that
$\sup_{v\in A}\Phi(v)\leq 0<\inf_{v\in B}\Phi(v)\quad\mbox{and}\quad\sup_{v\in X}\Phi(v)<c^{*}.$
Let us recall that $\ell=\delta q+b-N<\ell_{1},\ell_{2}$, where $\ell_{i}=\ell(\eta_{i},r_{i})=\delta r_{i}+\eta_{i}-N$. For all $\lambda\in\mathbb{R}$, $u\in\mathcal{M}$ and $t\geq 0$ we have
$\Phi(u_{t})=t^{\,\ell}\left(1-\frac{\lambda}{\tilde{\Psi}(u)}\right)-\frac{\mu t^{\,\ell_{2}}}{r_{2}}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{2}}}{|x|^{\eta_{2}}}dx-\frac{t^{\,\ell_{1}}}{2^{*}_{\eta_{1}}}\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx.$ (4.27)
Since $\mathcal{M}$ is bounded, there exist constants $c_{1},c_{2}>0$ such that
$\Phi(u_{t})\geq t^{\,\ell}\left(1-\frac{\lambda}{\tilde{\Psi}(u)}-c_{1}\mu t^{\,\ell_{2}-\ell}-c_{2}t^{\,\ell_{1}-\ell}\right)\geq\frac{t^{\,\ell}}{2}\left(1-\frac{\lambda}{\lambda_{k}}\right)>0,\quad\forall u\in B_{0},$
since $\tilde{\Psi}(u)\geq\lambda_{k}>\lambda$, provided $t=\rho$ is small enough. This gives us that
$\inf_{v\in B}\Phi(v)>0.$ (4.28)
Now, recall that $\lambda_{k+m-1}+\varepsilon<\lambda_{k+m}$ and $A_{0}\subset\mathcal{M}\backslash\tilde{\Psi}_{\lambda_{k+m-1}+\varepsilon}\subset\mathcal{M}\backslash\tilde{\Psi}_{\lambda_{k+m}}$. Then, from Lemma 4.5 we know that there exists $c_{3}>0$, independent on $\varepsilon$, such that
$\frac{1}{2^{*}_{\eta_{1}}}\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx\geq c_{3},\quad\forall u\in A_{0}.$
Since $\lambda,\mu\geq 0$, by (4.27) we obtain
$\Phi(u_{t})\leq t^{\,\ell}-\frac{t^{\,\ell_{1}}}{2^{*}_{\eta_{1}}}\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx\leq t^{\,\ell}\left(1-c_{3}t^{\,\ell_{1}-\ell}\right)\leq 0,\quad\forall u\in A_{0},$
for $t=R>\rho$, large enough. Hence,
$\sup_{v\in A}\Phi(v)\leq 0.$ (4.29)
It remains to estimate the supremum of $\Phi$ in $X$. Again by (4.27), since $\lambda,\mu\geq 0$, we get
$\Phi(u_{t})\leq t^{\,\ell}\left(1-\frac{\lambda}{\lambda_{k}+\varepsilon}\right)-c_{3}t^{\,\ell_{1}},\quad\forall u\in A_{0},\,t\geq 0.$
Thus,
$\sup_{u\in A_{0},0\leq t\leq R}\Phi(u_{t})$ $\leq$ $\max_{t\geq 0}\left[t^{\,\ell}\left(1-\frac{\lambda}{\lambda_{k}+\varepsilon}\right)-c_{3}t^{\,\ell_{1}}\right]=\frac{\ell_{1}-\ell}{\ell_{1}}\left(\frac{\ell}{c_{3}\ell_{1}}\right)^{\frac{\ell}{\ell_{1}-\ell}}\left(1-\frac{\lambda}{\lambda_{k}+\varepsilon}\right)^{\frac{\ell_{1}}{\ell_{1}-\ell}}$
$=$ $\frac{2-\eta_{1}}{N-\eta_{1}}\left(\frac{N-2}{c_{3}(N-\eta_{1})}\right)^{\frac{N-2}{2-\eta_{1}}}\left(1-\frac{\lambda}{\lambda_{k}+\varepsilon}\right)^{\frac{N-\eta_{1}}{2-\eta_{1}}}$
Choosing
$\delta_{k}=\lambda_{k}\left(c^{*}\frac{N-\eta_{1}}{2-\eta_{1}}\right)^{\frac{2-\eta_{1}}{N-\eta_{1}}}\left(\frac{c_{3}(N-\eta_{1})}{N-2}\right)^{\frac{N-2}{N-\eta_{1}}},$
we can see that for each $\lambda_{k}-\delta_{k}<\lambda<\lambda_{k}$, there exists $\varepsilon>0$ small enough such that
$\sup_{u\in A_{0},0\leq t\leq R}\Phi(u_{t})<c^{*}.$
Finally, combining this inequality, (4.28) and (4.29), the conclusion now follows by Proposition 4.6. $\Box$
In the next result $\lambda\in\mathbb{R}$ can be any number, even an eigenvalue for (1.1), and does not need to be close to an eigenvalue. However, we deal with subcritical perturbation $\mu|x|^{-\eta}|t|^{r_{2}-2}t$ and consider $\mu$ a large number.
Proofof Theorem 1.5. This proof is similar to the proof of previous theorem. Since $r_{2}<2^{*}_{\eta_{2}}$, by Lemma 4.3 we know that $\Phi$ satisfies the $(PS)_{c}$ for all $c<c^{*}(N,\eta_{1})$. Given $m\in\mathbb{N}$ we set
$k_{0}=\min\{k\in\mathbb{N}:\lambda<\lambda_{k}\}\quad\mbox{and}\quad\tilde{m}=k_{0}+m-1.$
Let $j\in\mathbb{N}$ and $\alpha\in\mathbb{R}$ be such that $\lambda_{\tilde{m}}=\cdots=\lambda_{\tilde{m}+j-1}<\alpha<\lambda_{\tilde{m}+j}$. Then, by Theorem 3.3(iii) we have
$i(\mathcal{M}\backslash\tilde{\Psi}_{\alpha})=\tilde{m}+j-1.$
Since $\mathcal{M}\backslash\tilde{\Psi}_{\alpha}$ is an open symmetric subset of $\mathcal{M}$, there exists a compact symmetric set $C\subset\mathcal{M}\backslash\tilde{\Psi}_{\alpha}$ with $i(C)=\tilde{m}+j-1$ (see the proof of Proposition 3.1 in Degiovanni and Lancelotti [11]). Set
$A_{0}=C\quad\mbox{and}\quad B_{0}=\tilde{\Psi}_{\lambda_{k_{0}}}.$
Hence, analyzing the cases $k_{0}=1$ or $k_{0}\geq 2$ for $B_{0}$, as done in the proof of Theorem 1.2, we see that
[…]
As before, for some $R>\rho>0$ to be chosen later, we set
$X=\{u_{t}:u\in A_{0},\,0\leq t\leq R\},\quad A=\{u_{R}:u\in A_{0}\}\quad\mbox{and}\quad B=\{u_{\rho}:u\in B_{0}\}.$
We claim that we may choose $R,\,\mu_{m}>0$ such that, for each $\mu\geq\mu_{m}$ there is $\rho\in(0,R)$ such that
$\sup_{v\in A}\Phi(v)\leq 0<\inf_{v\in B}\Phi(v)\quad\mbox{and}\quad\sup_{v\in X}\Phi(v)<c^{*}.$
Recalling (1.7), we have $\ell<\ell_{1},\ell_{2}$, where $\ell_{1}=\ell(\eta_{1},2^{*}_{\eta_{1}})$ and $\ell_{2}=\ell(\eta_{2},r_{2})$. For all $\lambda\in\mathbb{R}$, $u\in\mathcal{M}$ and $t\geq 0$ we have
$\Phi(u_{t})=t^{\,\ell}\left(1-\frac{\lambda}{\tilde{\Psi}(u)}\right)-\frac{\mu t^{\,\ell_{2}}}{r_{2}}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{2}}}{|x|^{\eta_{2}}}dx-\frac{t^{\,\ell_{1}}}{2^{*}_{\eta_{1}}}\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx.$ (4.30)
Due to the compactness of $A_{0}\subset\mathcal{M}$, there exists $c_{0}>0$ such that
$\frac{1}{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx,\quad\frac{1}{r_{2}}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{2}}}{|x|^{\eta_{2}}}dx,\quad\frac{1}{2^{*}_{\eta_{1}}}\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx\geq c_{0},\quad\forall u\in A_{0}.$
Denoting $\lambda^{-}=-\min\{\lambda,0\}$, by the boundedness of $\mathcal{M}$ in $L^{p}_{a}(\mathbb{R}^{N})$ we get
$1-\frac{\lambda}{\tilde{\Psi}(u)}\leq 1+\frac{\lambda^{-}}{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx\leq\beta,\quad\forall u\in A_{0},$
for some $\beta>0$. Since $\mu>0$, from (4.30) we obtain
$\Phi(u_{t})\leq\beta t^{\,\ell}-\frac{t^{\,\ell_{1}}}{2^{*}_{\eta_{1}}}\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx\leq\beta t^{\,\ell}-c_{0}t^{\,\ell_{1}}\leq 0,\quad\forall u\in A_{0},$
whether $t=R$, independent of $\mu$, is large enough. Hence
$\sup_{v\in A}\Phi(v)\leq 0.$ (4.31)
To estimate the supremum of $\Phi$ in $X$, using again (4.30) we obtain
$\sup_{u\in A_{0},0\leq t\leq R}\Phi(u_{t})$ $\leq$ $\max_{t\geq 0}\left[\beta t^{\,\ell}-\mu c_{0}t^{\,\ell_{2}}\right]=\frac{\ell_{2}-\ell}{\ell_{2}}\left(\frac{\beta^{\frac{\ell_{2}}{\ell}}\ell}{c_{0}\mu\ell_{2}}\right)^{\frac{\ell}{\ell_{2}-\ell}}.$
Picking $\mu_{m}>0$ sufficiently large, for $\mu\geq\mu_{m}$ it holds that
$\sup_{v\in X}\Phi(v)=\sup_{u\in A_{0},0\leq t\leq R}\Phi(u_{t})<c^{*}.$ (4.32)
Now we estimate $\Phi$ on $B$. By (4.30), due to the boundedness of $\mathcal{M}$, there exist $c_{1},c_{2}>0$ such that
$\Phi(u_{t})\geq t^{\,\ell}\left(1-\frac{\lambda^{+}}{\tilde{\Psi}(u)}-c_{1}\mu t^{\,\ell_{2}-\ell}-c_{2}t^{\,\ell_{1}-\ell}\right),$
where $\lambda^{+}=\max\{\lambda,0\}$. Hence, as $\tilde{\Psi}(u)\geq\lambda_{k_{0}}>\lambda^{+}$ for all $u\in B_{0}$ and $\ell_{1},\ell_{2}>\ell$, we can choose $t=\rho\in(0,R)$ small enough such that
$\inf_{u\in B_{0}}\Phi(u_{\rho})=\inf_{v\in B}\Phi(v)>0.$
Due to this last inequality, (4.31) and (4.32), the Proposition 4.6 can be applied to finish this proof. $\Box$
4.4 Problems with subscaled nonlinearities
We now focus on nonlinearities $f$ that display at least a subscaled pair $(r,\eta)$. For the proofs of Theorems 1.6-1.9, we make use of an abstract result in critical point theory (see [25, Proposition 3.36]), which we state here for reader’s convenience.
Proposition 4.8.
[…]
$\mathcal{F}_{k}=\{M\in\mathcal{F}:i(M)\geq k\}$
and set
$c_{k}:=\inf_{M\in\mathcal{F}_{k}}\sup_{u\in M}\Phi(u).$ (4.33)
If there exists a $k_{0}\geq 1$ such that $\infty<c_{k}<0$ for all $k\geq k_{0}$, then $c_{k_{0}}\leq c_{k_{0}+1}\leq\cdots\to 0$ is a sequence of critical values of $\Phi$.
Proofof Theorem 1.6. To proceed, let us recall the notations given in Definition 1 and (1.7). Since $(\eta,r)$ is a subscaled pair, we see that the potential operator $\tilde{f}$ induced by the nonlinearity satisfy
$\tilde{f}(u_{t})v_{t}=\int_{\mathbb{R}^{N}}\frac{|u_{t}|^{r-2}u_{t}v_{t}}{{|x|^{\eta}}}dx=o(t^{\,\ell})\|v\|\quad\mbox{as}\,\,t\to\infty,$
uniformly in $u$ on bounded sets of $E^{q}_{b}$, for all $v\in E^{q}_{b}$. Then, it is straightforward to verify that the associated functional, which in this case is given by
$\Phi(u)=\frac{1}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\frac{1}{q}\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx-\frac{1}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx,$
is coercive (see [23, Lemma 2.15] for a proof in a more general setting). Therefore, for any $c\in\mathbb{R}$, $(PS)_{c}$ sequences are always bounded. The compactness of the embedding $E^{q}_{b}\hookrightarrow L^{r}_{\eta}(\mathbb{R}^{N})$ implies that $\Phi$ satisfies the $(PS)_{c}$ conditions for all levels $c\in\mathbb{R}$, as a consequence of [23, Proposition 2.3].
The boundedness of $\Phi$} on bounded sets implies that $\Phi$} is bounded from below in $E^{q}_{b}$. Hence, the levels $c_{k}$ in (4.33) satisfy
$c_{k}\geq-C,\quad\mbox{for all}\,\,k\geq 1,$
for some positive constant $C$. Let us show that $c_{k}<0$ for all $k\geq 1$. For any $u\in\mathcal{M}$, we have
$\Phi(u_{t})$ $=$ $t^{\,\ell}-\frac{t^{\,\ell(\eta,r)}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx$
and since $\ell(\eta,r)<\ell$, this implies that
$\Phi(u_{t})$ $=$ $-t^{\,\ell(\eta,r)}\left(\frac{1}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx+o(1)\right)$ (4.34)
with $o(1)\to 0$ as $t\to 0^{+}$ uniformly in $u\in\mathcal{M}$. Now, for $k\geq 1$, in order to estimate the minimax level $c_{k}$ from above, let us set
$M_{t}=\{u_{t}:u\in\tilde{\Psi}^{\lambda_{k}}\},\quad\mbox{for}\quad t>0.$
We have $i(M_{t})=i(\tilde{\Psi}^{\lambda_{k}})$ as $u\mapsto u_{t}$ with $u\in\tilde{\Psi}^{\lambda_{k}}$, is an odd homeomorphism. Let $m\geq 1$ be such that $\lambda_{k}=\cdots=\lambda_{k+m-1}<\lambda_{k+m}$. By Theorem 3.3(iii) we have $i(\tilde{\Psi}^{\lambda_{k}})=k+m-1$. This implies that $i(M_{t})=k+m-1\geq k,$ and so $M_{t}\in\mathcal{F}_{k}$. Using (4.34) and Lemma 4.5 we see that
$\Phi(u_{t})\leq-t^{\,\ell(\eta,r)}\left(c+o(1)\right)<0\quad\mbox{for all}\quad u\in\tilde{\Psi}^{\lambda_{k}}$
for $t>0$ sufficiently small. Then
$c_{k}=\inf_{M\in\mathcal{F}_{k}}\sup_{u\in M}\Phi(u)\leq\sup_{u\in M_{t}}\Phi(u)=\sup_{u\in\tilde{\Psi}^{\lambda_{k}}}\Phi(u_{t})<0,$
for all $k\geq 1$. By Proposition 4.8, this concludes the proof. $\Box$
Proofof Theorem 1.7. The associated functional in this case is given by
$\Phi(u)=\frac{1}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\frac{1}{q}\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx-\frac{\lambda}{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx-\frac{1}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx+\frac{1}{r_{1}}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{1}}}{|x|^{\eta_{1}}}dx.$
Then, if $\lambda\leq 0$ we get
$\Phi(u)\geq\frac{1}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\frac{1}{q}\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx-\frac{1}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx,$
[…]
$\frac{\lambda}{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}\,dx\;\leq\;\frac{1}{r_{1}}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{1}}}{|x|^{\eta_{1}}}\,dx\;+\;\tilde{C}\int_{\mathbb{R}^{N}}\frac{|u|^{\tilde{r}}}{|x|^{\tilde{\eta}}}\,dx.$
This gives us
$\Phi(u)\geq\tilde{\Phi}(u):=\frac{1}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\frac{1}{q}\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx-\frac{1}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx-\tilde{C}\int_{\mathbb{R}^{N}}\frac{|u|^{\tilde{r}}}{|x|^{\tilde{\eta}}}\,dx.$
[…]
So, let us see the case $r_{1}=2^{*}_{\eta_{1}}$. Let $(u_{n})$ be a $(PS)_{c}$ sequence, for some $c\in\mathbb{R}$. Since $(u_{n})$ is bounded, let $u\in E^{q}_{b}$ be its weak limit, up to a subsequence. Then, we follow the same steps of the proof of Lemma 4.3 until (4.13). In this case here, we get
$\int_{\mathbb{R}^{N}}|\nabla(u_{n}-u)|^{2}dx+\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{q}}{|x|^{b}}dx=-\int_{\mathbb{R}^{N}}\frac{|u_{n}-u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx+o(1).$
This gives that $(u_{n})$ converges strongly to $u$ both in $E^{q}_{b}$ and $L^{2^{*}_{\eta_{1}}}_{\eta_{1}}(\mathbb{R}^{N})$. Since $\Phi$} is bounded from below in $E^{q}_{b}$, as in the proof of the previous theorem, the levels $c_{k}$ in (4.33) are bounded from below. We just need to check that $c_{k}<0$ for all $k\geq 1$. Notice that now we have, for any $u\in\mathcal{M}$,
$\Phi(u_{t})$ $=$ $t^{\,\ell}-\frac{t^{\,\ell(\eta,r)}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx-\frac{\lambda t^{\,\ell}}{p}\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx+\frac{t^{\,\ell(\eta_{1},r_{1})}}{r_{1}}\int_{\mathbb{R}^{N}}\frac{|u|^{r_{1}}}{|x|^{\eta_{1}}}dx$
and since $\ell(\eta,r)<\ell$ and $\ell(\eta,r)<\ell(\eta_{1},r_{1})$, this implies that
$\Phi(u_{t})$ $=$ $-t^{\,\ell(\eta,r)}\left(\frac{1}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx+o(1)\right)$
with $o(1)\to 0$ as $t\to 0^{+}$ uniformly in $u\in\mathcal{M}$. This is exactly the same identity as in (4.34), and therefore the remainder of the argument coincides with the proof of Theorem 1.6. $\Box$
Proofof Theorem 1.8. The associated functional, given in (1.6), is in this setting given by
$\Phi(u)=\Phi_{\lambda}(u)-\frac{1}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx.$
As in the proof of Theorem 1.6, the potential operator $\tilde{f}$ satisfy
$\tilde{f}(u_{t})v_{t}=\int_{\mathbb{R}^{N}}\frac{|u_{t}|^{r-2}u_{t}v_{t}}{{|x|^{\eta}}}dx=o(t^{\,\ell})\|v\|\quad\mbox{as}\,\,t\to\infty,$
uniformly in $u$ on bounded sets of $E^{q}_{b}$, for all $v\in E^{q}_{b}$. So, the $(PS)_{c}$ condition holds for all levels $c\in\mathbb{R}$ provided $\lambda$ is not an eigenvalue. This is proved in a more general framework in [23, Lemma 2.26]. We also note that it suffices to consider the case $\lambda>0$, since when $\lambda\leq 0$ the functional is coercive and the argument follows exactly as in the proof of Theorem 1.6. Moreover, if we consider $c_{k}$ as in (4.33), the estimate $c_{k}<0$ is also exactly the same, since here we have, for all $u\in\mathcal{M}$,
$\Phi(u_{t})$ $=$ $t^{\,\ell}\left(1-\frac{\lambda}{\tilde{\Psi}(u)}\right)-\frac{t^{\,\ell(\eta,r)}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx=-t^{\,\ell(\eta,r)}\left(\frac{1}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx+o(1)\right)$ (4.35)
with $o(1)\to 0$ as $t\to 0^{+}$ uniformly in $u\in\mathcal{M}$.
The main difference is that we need to estimate $c_{k}$ for some $k\geq k_{0}\geq 1$ from below. Coercivity is no longer available since we are assuming $\lambda>0$. Note that, as $\lambda$ is not an eigenvalue, we may have $0<\lambda<\lambda_{1},$ or $\lambda_{k-1}<\lambda<\lambda_{k}$ for some $k\geq 2$. Set $k_{0}:=\min\{k\in\mathbb{N}:\lambda<\lambda_{k}\}.$ By the definition of $\mathcal{F}_{k}$ in Proposition 4.8, for each $M\in\mathcal{F}_{k}$ we have
$i(M)\geq k\geq k_{0},$ (4.36)
for all $k\geq k_{0}\geq 1$. Setting $Y=\{u_{t}:u\in\tilde{\Psi}_{\lambda_{k_{0}}},\,t\geq 0\},$ we claim that
$M\cap Y\neq\emptyset\quad\mbox{for all}\,\,M\in\mathcal{F}_{k}\,\,\mbox{and}\,\,k\geq k_{0}.$ (4.37)
In fact, if this were not true, we would have $M\cap Y=\emptyset$ for some $M\in\mathcal{F}_{k}$ and $k\geq k_{0}$. So $\pi_{|_{M}}:M\to\mathcal{M}\backslash\tilde{\Psi}_{\lambda_{k_{0}}}$ would be an odd continuous map, and hence we would have $i(M)\leq i(\mathcal{M}\backslash\tilde{\Psi}_{\lambda_{k_{0}}}).$ On the other hand, if $\lambda<\lambda_{1}$, then $k_{0}=1$ and by Theorem 3.3(i) we have $\tilde{\Psi}_{\lambda_{1}}=\mathcal{M}$ so that $i(\mathcal{M}\backslash\tilde{\Psi}_{\lambda_{1}})=i(\emptyset)=0,$ whereas if $\lambda_{k_{0}-1}<\lambda<\lambda_{k_{0}}$ with $k_{0}\geq 2$, the same Theorem 3.3(iii) yields $i(\mathcal{M}\backslash\tilde{\Psi}_{\lambda_{k_{0}}})=k_{0}-1.$ In both cases, we would conclude that
$i(M)\leq i(\mathcal{M}\backslash\tilde{\Psi}_{\lambda_{k_{0}}})\leq k_{0}-1,$
which contradicts (4.36). Thus, (4.37) holds. Therefore, for each $M\in\mathcal{F}_{k}$}
$\sup_{u\in M}\Phi(u)\geq\sup_{u\in M\cap Y}\Phi(u)\geq\inf_{u\in M\cap Y}\Phi(u)\geq\inf_{u\in Y}\Phi(u).$
Since $\tilde{\Psi}(u)\geq\lambda_{k_{0}}>\lambda$ for all $u\in\tilde{\Psi}_{\lambda_{k_{0}}}$ and $\mathcal{M}$ is bounded in $L^{r}_{\eta}(\mathbb{R}^{N})$ and in $L^{p}_{a}(\mathbb{R}^{N})$, the first equality in (4.35) yields $\inf_{u\in Y}\Phi(u):=-C>$-$\infty$. Hence
$c_{k}=\inf_{M\in\mathcal{F}_{k}}\sup_{u\in M}\Phi(u)\geq\inf_{u\in Y}\Phi(u)\geq-C,\quad\forall k\geq k_{0},$
and by Proposition 4.8 we may conclude the proof. $\Box$
Proofof Theorem 1.9. This proof also applies Proposition 4.8. Here, as in [23], we use a truncation of the functional $\Phi_{\mu}$ associated with this problem, which is
$\Phi_{\mu}(u)=\frac{1}{2}\int_{\mathbb{R}^{N}}|\nabla u|^{2}dx+\frac{1}{q}\int_{\mathbb{R}^{N}}\frac{|u|^{q}}{|x|^{b}}dx-\frac{\mu}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx-\frac{1}{2^{*}_{\eta_{1}}}\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx.$
We verify that for any admissible pair $(\eta,r)$, since $r\geq 2^{*}_{b,q,\eta}$, there exists $C>0$ such that
$\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx\leq C[I(u)]^{\frac{\ell(\eta,r)}{\ell}},\quad\forall u\in E^{q}_{b}.$ (4.38)
Then
$\Phi_{\mu}(u)$ $\geq$ $I(u)-{C}\mu[I(u)]^{\frac{\ell(\eta,r)}{\ell}}-{C}_{1}[I(u)]^{\frac{\ell(\eta_{1},2^{*}_{\eta_{1}})}{\ell}}\quad\forall u\in E^{q}_{b}.$
Let us denote $\gamma_{\mu}(t)=t-{C}\mu t^{\frac{\ell(\eta,r)}{\ell}}-{C}_{1}t^{\frac{\ell(\eta_{1},2^{*}_{\eta_{1}})}{\ell}}$, for $t\geq 0$. So, we can write
$\Phi_{\mu}(u)$ $\geq$ $\gamma_{\mu}(I(u)),\quad\forall u\in E^{q}_{b}.$
Once $\ell(\eta_{1},2^{*}_{\eta_{1}})>\ell>\ell(\eta,r)>0$ there is $\mu^{*}>0$ such that for each $\mu\in(0,\mu^{*})$ it is possible to find $R_{1}(\mu)$ and $R_{2}(\mu)>0$ such that
$\gamma_{\mu}(t)<0,\,\,\forall t\in[0,R_{1}(\mu))\cup(R_{2}(\mu),\infty)\qquad\mbox{and}\qquad\gamma_{\mu}(t)\geq 0,\,\,\forall t\in[R_{1}(\mu)),(R_{2}(\mu)].$
For $\mu\in(0,\mu^{*})$ it holds that $\gamma_{\mu}(t)\geq\gamma_{\mu^{*}}(t)$ for all $t\in[R_{1}(\mu^{*})),(R_{2}(\mu^{*})]$, hence, $R_{1}(\mu)\leq R_{1}(\mu^{*})$ and $R_{2}(\mu^{*})\leq R_{2}(\mu)$. Consider now $\xi_{\mu}:[0,\infty)\to[0,1]$ smooth and satisfying
$\xi_{\mu}\equiv 1\,\,\mbox{in}\,\,[0,R_{1}(\mu)]\quad\mbox{and}\quad\xi_{\mu}\equiv 0\,\,\mbox{in}\,\,[R_{2}(\mu),\infty),$
and define the truncated functional $\tilde{\Phi}_{\mu}:E^{q}_{b}\to\mathbb{R}$ by
$\tilde{\Phi}_{\mu}(u)=\xi_{\mu}(I(u)){\Phi}_{\mu}(u).$
Note that if $\tilde{\Phi}_{\mu}(u)<0$ then $I(u)\in(0,R_{1}(\mu)),$ and, by the continuity of $I$, $I(v)\in(0,R_{1}(\mu))$ for $v$ in some neighborhood of $u$ in $E^{q}_{b}$, which imply
$\tilde{\Phi}_{\mu}(u)={\Phi}_{\mu}(u)<0\quad\mbox{and}\quad\tilde{\Phi}^{\prime}_{\mu}(u)={\Phi}^{\prime}_{\mu}(u).$ (4.39)
Hence critical points for $\tilde{\Phi}_{\mu}$ at negative levels are actual critical points for ${\Phi}_{\mu}$. Thus, to conclude this proof is enough to apply Proposition 4.8 to find a sequence of critical points for $\tilde{\Phi}_{\mu}$ whose energy is negative. Our first step is to show that $\tilde{\Phi}_{\mu}$ satisfies the $(PS)_{c}$ condition for $c<0$, provided $\mu\in(0,\mu^{*})$ for some $\mu^{*}$ sufficiently small. Let us consider then a sequence $(u_{n})$ in $E^{q}_{b}$ such that
$\tilde{\Phi}_{\mu}(u_{n})\to c<0\quad\text{and}\quad\tilde{\Phi}^{\prime}_{\mu}(u_{n})\to 0.$
Since $\tilde{\Phi}_{\mu}(u_{n})<0$ for large $n$, we have $I(u_{n})\in(0,R_{1}(\mu^{*}))$, which forces $(u_{n})$ to be uniformly bounded for $\mu\in(0,\mu^{*})$. Up to a subsequence, we may assume $u_{n}\rightharpoonup u$ in $E^{q}_{b}$, and this implies $I(u)\leq R_{1}(\mu^{*})$, for $\mu\in(0,\mu^{*})$. Then, (4.38) ensures the uniform boundedness of $\|u\|_{L^{r}_{\eta}}$. Note that (4.39) also implies that $(u_{n})$ is a $(PS)_{c}$ sequence for ${\Phi}_{\mu}$. Arguing as in Lemma 4.3, if $(u_{n})$ has no subsequence convergent in $E^{q}_{b}$, we see that
$c$ $\geq$ $\frac{2-\eta_{1}}{2(N-\eta_{1})}S_{\eta_{1}}^{\frac{N-\eta_{1}}{2-\eta_{1}}}-\mu\tilde{d}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx$
for some positive constant $\tilde{d}$ independent of $u$ and $\mu$ (see (4.17) and recall that here $\ell(\eta,r)<\ell$). Hence, the uniform boundedness of $\|u\|_{L^{r}_{\eta}}$ for $\mu\in(0,\mu^{*})$ and the fact that $c<0$ gives us
$\mu C\tilde{d}\geq\mu\tilde{d}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx\geq\frac{2-\eta_{1}}{2(N-\eta_{1})}S_{\eta_{1}}^{\frac{N-\eta_{1}}{2-\eta_{1}}}>0$
which is not possible if $\mu^{*}>0$ sufficiently small. By considering $\mu^{*}>0$ smaller than the first choice, if necessary, we finally conclude that $\tilde{\Phi}_{\mu}$ satisfies the $(PS)_{c}$ condition for any $c<0$ and $\mu\in(0,\mu^{*})$. Now, let us pick $\mu\in(0,\mu^{*})$ and estimate the levels $c_{k}$ defined in (4.33). Since $\tilde{\Phi}_{\mu}$ is bounded, we have $c_{k}>$-$\infty$ for all $k\in\mathbb{N}$. On the other hand, since $I(u_{t})=t^{\,\ell}<R_{1}(\mu)$ for all $u\in\mathcal{M}$, if $t>0$ is small, this gives us $\tilde{\Phi}_{\mu}(u_{t})=\Phi_{\mu}(u_{t})$. Recalling that $\ell(\eta_{1},2^{*}_{\eta_{1}})>\ell>\ell(\eta,r)>0$, for all $u\in\mathcal{M}$ we get
$\Phi_{\mu}(u_{t})$ $=$ $t^{\ell}-\frac{\mu t^{\,\ell(\eta,r)}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx-\frac{t^{\,\ell(\eta_{1},2^{*}_{\eta_{1}})}}{2^{*}_{\eta_{1}}}\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx=-t^{\,\ell(\eta,r)}\left(\frac{\mu}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx+o(1)\right),$
with $o(1)\to 0$ as $t\to 0^{+}$ uniformly in $u\in\mathcal{M}$. As in the proof of Theorem 1.6 (see (4.34) and below), for $M_{t}=\{u_{t}:u\in\tilde{\Psi}^{\lambda_{k}}\}$ we see that $M_{t}\in\mathcal{F}_{k}$ and, by Lemma 4.5, we obtain $\sup_{u\in\tilde{\Psi}^{\lambda_{k}}}\Phi(u_{t})<0$ for $t>0$ sufficiently small. Hence
$c_{k}=\inf_{M\in\mathcal{F}_{k}}\sup_{u\in M}\tilde{\Phi}_{\mu}(u)\leq\sup_{u\in M_{t}}\tilde{\Phi}_{\mu}(u)=\sup_{u\in\tilde{\Psi}^{\lambda_{k}}}\Phi_{\mu}(u_{t})<0,$
and by Proposition 4.8, this concludes the proof. $\Box$
We now arrive at the proof of the final main theorem of this work.
Proofof Theorem 1.10. First, notice that the associated functional satisfies the $(PS)_{c}$ condition for all $c<c^{*}:=\frac{2-\eta_{1}}{2(N-\eta_{1})}S_{\eta_{1}}^{\frac{N-\eta_{1}}{2-\eta_{1}}}$, as observed in Remark 4.2. Then, we shall apply Corollary 4.7. For any $\lambda\in(\lambda_{k},\infty)$ we may find $m\in\mathbb{N}$ such that $\lambda_{k+m-1}<\lambda\leq\lambda_{k+m}$. We fix $\tilde{\lambda}\in(\lambda_{k+m-1},\lambda)$ and $\lambda^{\prime}\in(\lambda,\infty)$. By Theorem 3.3(iii) we have
$i(\mathcal{M}\backslash\tilde{\Psi}_{\tilde{\lambda}})=k+m-1\geq k.$
Pick $A_{0}\subset\mathcal{M}\backslash\tilde{\Psi}_{\tilde{\lambda}}$ a compact symmetric subset of index k+m1k+m-1 (see the proof of Proposition 3.1 in [11]). For any $R>\rho>0$ to be chosen later, we set
$A=\{u_{R}:u\in A_{0}\}\quad\mbox{and}\quad X=\{u_{t}:u\in A_{0},\,t\geq 0\}.$
For all $u\in\mathcal{M}$ we have
$\Phi(u_{t})=t^{\,\ell}\left(1-\frac{\lambda}{\tilde{\Psi}(u)}\right)+\frac{\mu\,t^{\,\ell(\eta,r)}}{r}\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx-\frac{t^{\,\ell(\eta_{1},2^{*}_{\eta_{1}})}}{2^{*}_{\eta_{1}}}\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}dx.$ (4.40)
Note that as $\lambda^{\prime}>\lambda>\lambda_{1}$, the boundedness of $\|u\|_{L^{2^{*}_{s}}}$ for $u\in\mathcal{M}$, Theorem 3.3(i) and Lemma 4.5 give us the estimate
$\Phi(u_{t})\geq-t^{\,\ell}\left(\frac{\lambda}{\lambda_{1}}-1\right)+c_{1}\mu t^{\,\ell(\eta,r)}-c_{2}t^{\,\ell(\eta_{1},2^{*}_{\eta_{1}})}\quad\forall u\in\tilde{\Psi}^{\lambda^{\prime}},\forall t\geq 0,$
for some $c_{1},c_{2}>0$ which do not depend on $\mu$. Also observe that
$\Phi(u_{t})\geq t^{\,\ell}\left(1-\frac{\lambda}{\lambda^{\prime}}\right)-c_{2}t^{\,\ell(\eta_{1},2^{*}_{\eta_{1}})}\quad\forall u\in\mathcal{M}\backslash\tilde{\Psi}^{\lambda^{\prime}},\forall t\geq 0.$
Recalling that $\ell(\eta_{1},2^{*}_{\eta_{1}})>\ell>\ell(\eta,r)$ we conclude that
$\inf_{u\in\mathcal{M}}\Phi(u_{\rho})>0$
provided $\rho>0$ is small enough. On the other hand, since $A_{0}\subset\mathcal{M}\backslash\tilde{\Psi}_{\tilde{\lambda}}$ we have $\tilde{\Psi}(u)<\tilde{\lambda}$ for any $u\in A_{0}$. Moreover, $\inf_{u\in A_{0}}\|u\|_{L^{2^{*}_{\eta_{1}}}_{\eta_{1}}}>0$, as $A_{0}$ is compact. Hence, using the boundedness of $\|u\|_{L^{r}_{\eta}}$ for $u\in\mathcal{M}$, by (4.40) we get
$\Phi(u_{t})\leq-t^{\,\ell}\left(\frac{\lambda}{\tilde{\lambda}}-1\right)+c^{\prime}\mu t^{\,\ell(\eta,r)}-c_{3}t^{\,\ell(\eta_{1},2^{*}_{\eta_{1}})}\quad\forall u\in A_{0},\forall t\geq 0,$
for some $c^{\prime},c_{3}>0$. Picking now $R>\rho$ sufficiently large, we obtain
$\sup_{u\in A}\Phi(u)=\sup_{u\in A_{0}}\Phi(u_{R})\leq 0.$
Note also that, for $u\in A_{0}$ and $t\geq 0$, we have
$\Phi(u_{t})\leq$ $\ c^{\prime}\mu t^{\,\ell(\eta,r)}-c_{3}t^{\,\ell(\eta_{1},2^{*}_{\eta_{1}})}$
$\leq$ $\ \big(\ell(\eta_{1},2^{*}_{\eta_{1}})-\ell(\eta,r)\big)\left[\frac{\ell(\eta,r)}{c_{3}}\right]^{\frac{\ell(\eta,r)}{\ell(\eta_{1},2^{*}_{\eta_{1}})-\ell(\eta,r)}}\left[\frac{c^{\prime}\mu}{\ell(\eta_{1},2^{*}_{\eta_{1}})}\right]^{\frac{\ell(\eta_{1},2^{*}_{\eta_{1}})}{\ell(\eta_{1},2^{*}_{\eta_{1}})-\ell(\eta,r)}}.$
As a consequence, for $\mu\in(0,\mu^{*})$ with $\mu^{*}$ sufficiently small, it holds that
$\sup_{u\in X}\Phi(u)=\sup_{u\in A_{0},t\in[0,R]}\Phi(u_{t})<\frac{2-\eta_{1}}{2(N-\eta_{1})}S_{\eta_{1}}^{\frac{N-\eta_{1}}{2-\eta_{1}}},$
and therefore, by Corollary 4.7 there exist $k+m-1(\geq k)$ pairs of critical points for $\Phi$, at positive energy levels. This concludes the proof. $\Box$
5 Radial improvements
This section is devoted to exploring the improvements that arise when working in the radial subspace $E^{q}_{b,\mathrm{rad}}\subset E^{q}_{b}$. We begin by showing that the embeddings in Lemma 2.2 admit a strictly larger range of exponents in the radial setting when $0\leq\eta<b$. In contrast, when $\eta\geq b$, no improvement is obtained, and the embedding ranges coincide with those of the nonradial case. In particular, for $N=2$ we prove that one can work with singular weights $|x|^{-\eta}$ even when $\eta<b$, a feature that is not available in the nonradial setting because of the restrictions in Lemma 2.2. The main tool leading to these improved ranges is provided by [22, Theorem 18 and Proposition 26], where refined versions of the Caffarelli–Kohn–Nirenberg inequalities are established for radial smooth functions. After proving the radial embeddings, we list some existence results for radial solutions of (1.4) and compare them with those obtained in the nonradial case.
For $N\geq 2$ and $0\leq\eta<b<2<q$, we denote
$2^{\text{rad}}_{b,q,\eta}:=\frac{q(2N-2-\eta)+2(b-\eta)}{2N-2-b}.$
Lemma 5.1.
Suppose $N\geq 2$, let $0<b<2<q<2^{*}_{b}$ and $0\leq\eta<b$. Then we have $E^{q}_{b,rad}\hookrightarrow L^{r}_{\eta}(\mathbb{R}^{N})$ for
$\begin{cases}\vskip 5.69046pt\par r\in\left[2^{\rm{rad}}_{b,q,\eta},\infty\right)&\text{if}\quad N=2,\\ \vskip 5.69046pt\par r\in\left[2^{\rm{rad}}_{b,q,\eta},2^{*}_{\eta}\right]&\text{if}\quad N\geq 3,\end{cases}$
and for $\theta\in\left(0,1\right]$ such that $\frac{1}{r}=\theta\frac{N-2}{2(N-\eta)}+(1-\theta)\frac{N-b}{q(N-\eta)}$ it holds
$\left(\int_{\mathbb{R}^{N}}\frac{|u|^{r}}{|x|^{\eta}}dx\right)^{\frac{1}{r}}\leq C\|\nabla u\|^{\theta}_{2}\||\cdot|^{\frac{-b}{q}}u\|_{q}^{1-\theta},\quad\forall u\in E^{q}_{b,rad}(\mathbb{R}^{N}).$
Moreover, all these embeddings are compact except for the endpoints of these intervals.
Proof.
We are going to apply CKN inequality again, for $\gamma=-\eta/r$. From the proof of Lemma 2.2, we know that when $N\geq 3$, $\sigma=-\eta/(r\theta)+(1-\theta)b/(q\theta)\leq 0$ for all $\theta\in[\theta_{1},1]$ and $\sigma\geq-1$ if $\theta=1$, since $\eta\leq 2$. Then (2.2) holds for any $u\in C_{c}^{\infty}(\mathbb{R}^{N})$, in particular for radial functions.
On the other hand, for $\theta\in(0,\theta_{1})$ if $N\geq 3$, or $\theta\in(0,1)$ if $N=2$, we have $\sigma>0$ so that (2.4) is not satisfied. Nevertheless, for radial functions we can apply [22, Theorem 18] which guarantees that (2.2) still holds if we assume (2.1) with $\theta>0$, (2.3) and consider $\sigma\in(0,N-1]$. We verify that, for $\theta\in(0,1)$, one has
$N-1\geq\sigma=-\frac{\eta(N-2)}{2(N-\eta)}-\frac{(1-\theta)N(\eta-b)}{\theta q(N-\eta)}\quad\Longleftrightarrow\quad\theta\geq\frac{2(b-\eta)}{2(b-\eta)+q(2N-2-\eta)}=:\theta_{2}.$
Moreover, for $\theta=\theta_{2}$ we get $r=2^{\text{rad}}_{b,q,\eta}$.
To obtain compact embeddings away from the endpoints, one argues exactly as in the proof of Lemma 2.2, but using [22, Proposition 26] instead of Proposition 2.1, which allows $\sigma<N-1$. This completes the proof. ∎
Remark 5.1.
Notice that if $\eta<b$ and $N\geq 3$, then from the definitions of $2^{*}_{b,q,\eta}$ and $2^{\mathrm{rad}}_{b,q,\eta}$ we obtain
$\frac{q(N-\eta)}{N-b}\;<\;2^{\mathrm{rad}}_{b,q,\eta}\;<\;2^{*}_{b,q,\eta},$
so the admissible range of exponents is strictly larger in the radial setting. Furthermore, the exponent $2^{\mathrm{rad}}_{b,q,\eta}$ is well defined when $N=2$ and $\eta<b$, whereas $2^{*}_{b,q,\eta}$ is not defined in this regime. Consequently, for $N=2$ and $\eta<b$ we obtain an entirely new interval of Sobolev exponents, allowing us to treat the problem (1.4) in this setting.
Remark 5.2.
This last lemma becomes particularly interesting when considering the case $\eta=0$, that is, embeddings of $E^{q}_{b}$ into the non-weighted Lebesgue spaces $L^{r}$. Notice that, in the nonradial case (Lemma 2.2), we cannot take $\eta=0$ when $N=2$, and if $N\geq 3$ the admissible interval reduces to the single point $r=2^{*}$. As a consequence, no compact embeddings are available. The radial case presents a better situation: in this setting, there exist nontrivial intervals for $r$ in any dimension $N\geq 2$, and the embeddings are compact except at the endpoints.
All the existence and multiplicity results stated in Theorems 1.11.10, which were formulated under the admissibility condition of Definition 1 and rely only on the weighted Sobolev embeddings from Lemma 2.2, remain valid, with exactly the same proofs, when restricted to the radial subspace $E^{q}_{b,\mathrm{rad}}$, provided that the admissibility requirement on $(\eta,r)$ is replaced by the corresponding radial embedding ranges from Lemma 5.1.
For $\eta\geq b$, the radial and nonradial admissible intervals coincide, so the statements of all theorems remain unchanged. When $0\leq\eta<b$ and $N\geq 3$, however, the radial lower endpoint improves from $2^{*}_{b,q,\eta}$ to $2^{\mathrm{rad}}_{b,q,\eta}$, thereby enlarging the class of superscaled admissible pairs. Consequently, the hypotheses of Theorems 1.11.5 and 1.7 may be relaxed in the radial setting. The remaining results either involve subscaled pairs, which never occur when $\eta<b$, or require critical exponents, and therefore do not benefit from the radial improvement.
Finally, when $N=2$ and $0\leq\eta<b$, Lemma 5.1 provides a genuinely new admissible interval [ $2^{\mathrm{rad}}_{b,q,\eta},\infty)$ for $r$, provided $(\eta,r)$ is superscaled. Thus, the results in Theorems 1.1, 1.2 and 1.7 admit radial analogues even in two dimensions, where such pairs did not exist in the nonradial case.
6 Appendix
In this section, we prove the density of smooth functions in the functional spaces that make up the framework of this work.
Proposition 6.1.
Consider $N\geq 2$, $b\in(0,N)$ and $q\geq 1$. If $N=2$ assume also that $q\geq b$. Then $C^{\infty}_{c}(\mathbb{R}^{N})$ is dense in $E^{q}_{b}$.
Proof.
Consider $\varphi\in C^{\infty}_{c}(\mathbb{R}^{N})$ such that $0\leq\varphi(x)\leq 1$, $\|\nabla\varphi\|_{\infty}\leq 2$, $\varphi\equiv 1$ in $B_{1}$ and $\varphi\equiv 0$ in $B_{2}^{c}$. Then define the sequence $\varphi_{k}(x)=\varphi(x/k)$. Consider also the sequence of functions $T_{k}:\mathbb{R}\to\mathbb{R}$ given by
$T_{k}(s)=\begin{cases}s&\text{if}\quad|s|<k\\ k\frac{s}{|s|}&\text{if}\quad|s|\geq k.\end{cases}$
For a fixed $u\in E^{q}_{b}$ we denote $u_{k}:=\varphi_{k}T_{k}(u)$. It is easily seen that $u_{k}\to u$ in $L^{q}_{b}(\mathbb{R}^{N})$. Moreover,
$\left\|\frac{\partial u_{k}}{\partial x_{i}}-\frac{\partial u}{\partial x_{i}}\right\|_{L^{2}}\leq\left\|\frac{\partial\varphi_{k}}{\partial x_{i}}T_{k}(u)\right\|_{L^{2}}+\left\|\left(\frac{\partial(T_{k}(u))}{\partial x_{i}}-\frac{\partial u}{\partial x_{i}}\right)\varphi_{k}\right\|_{L^{2}}+\left\|\frac{\partial u}{\partial x_{i}}(1-\varphi_{k})\right\|_{L^{2}}.$ (6.1)
Since $\varphi_{k}(x)\to 1$ and $\frac{\partial(T_{k}(u))}{\partial x_{i}}(x)\to\frac{\partial u}{\partial x_{i}}(x)$ a.e. in $\mathbb{R}^{N}$ and $\frac{\partial u}{\partial x_{i}}\in L^{2}(\mathbb{R}^{N})$, these last two norms go to 0 as $k\to\infty$. Moreover, for $i\in\{1,\cdots,N\}$, it holds
$\int_{\mathbb{R}^{N}}\left|\frac{\partial\varphi_{k}}{\partial x_{i}}T_{k}(u)\right|^{2}dx=\frac{1}{k^{2}}\int_{\mathbb{R}^{N}}\left|\frac{\partial\varphi}{\partial x_{i}}(x/k)T_{k}(u(x))\right|^{2}dx.$
Here we analyze the cases $N=2$ or grater. When $N\geq 3$ we have $E^{q}_{b}\hookrightarrow L^{2^{*}}(\mathbb{R}^{N})$. Then, using $|T_{k}(s)|\leq|s|$ and $supp(\frac{\partial\varphi_{k}}{\partial x_{i}})\subset B_{2k}\setminus B_{k}$, and the Holder’s inequality we get
$\int_{\mathbb{R}^{N}}\left|\frac{\partial\varphi_{k}}{\partial x_{i}}T_{k}(u)\right|^{2}dx\leq\|\nabla\varphi\|_{L^{N}}^{2}\left(\int_{B_{2k}\setminus B_{k}}|u|^{2^{*}}dx\right)^{\frac{2}{2^{*}}}=o(1),\quad\text{as}\,\,k\to\infty.$
Assume now $N=2$ and $b\in(0,2)$. If $\max\{b,1\}\leq q\leq 2$, since $|T_{k}(s)|\leq\min\{|s|,k\}$ and $u\in L^{q}_{b}(\mathbb{R}^{2})$, we have
$\int_{\mathbb{R}^{2}}\left|\frac{\partial\varphi_{k}}{\partial x_{i}}T_{k}(u)\right|^{2}dx$ $\leq\frac{k^{2-q}}{k^{2}}\|\nabla\varphi\|_{\infty}^{2}\int_{B_{2k}\setminus B_{k}}|u(x)|^{q}dx\leq\frac{2^{b+2}}{k^{q-b}}\int_{B_{2k}\setminus B_{k}}\frac{|u(x)|^{q}}{|x|^{b}}dx=o(1).$
In case $q>2$, using Holder’s inequality again we get
$\int_{\mathbb{R}^{2}}\left|\frac{\partial\varphi_{k}}{\partial x_{i}}T_{k}(u)\right|^{2}dx$ $\leq k^{-\frac{2(2-b)}{q}}\|\nabla\varphi\|_{L^{\frac{2q}{q-2}}}^{2}\|u\|_{L^{q}_{b}}^{2}=o(1).$
Thus, for $N=2$, $b\in(0,2)$ and $q\geq\max\{b,1\}$ or $N\geq 3$, $b\in(0,N)$ and $q\geq 1$, from (6.1) we obtain $\nabla u_{k}\to\nabla u$ in $(L^{2}(\mathbb{R}^{N}))^{N}$. Therefore, $u_{k}\to u$ in $E^{q}_{b}$.
Now, let us deal with smooth functions with compact support. Given $\varepsilon>0$, fix $k_{0}$ such that $u_{0}:=u_{k_{0}}$ satisfies
$\|u_{0}-u\|<\frac{\varepsilon}{2}.$ (6.2)
Since $b\in(0,N)$ we have $w(x):=|x|^{-b}$ a function in $L^{1}_{loc}(\mathbb{R}^{N})$. So, $w\chi_{\{w\leq n^{q}\}}\to w$ in $L^{1}(B_{2k_{0}})$ as $n\to\infty$ and we fix $n_{0}\in\mathbb{N}$ such that
$\|w\chi_{\{w\leq n^{q}_{0}\}}-w\|_{L^{1}(B_{2k_{0}})}<\frac{\varepsilon}{10k_{0}}.$ (6.3)
Also, choose $\sigma>0$ such that
$\|w\|_{L^{1}(B_{2k_{0}+\sigma}\setminus B_{2k_{0}})}<\frac{\varepsilon}{10k_{0}}.$ (6.4)
At this point, we fix a function $\rho\in C^{\infty}_{c}(\mathbb{R}^{N})$ such that $\rho\geq 0$, $supp(\rho)\subset B_{1}$ and $\|\rho\|_{L^{1}}=1$, and then define $\rho_{m}(x)=m^{N}\rho(mx)$. We know that $\rho_{m}*v\to v$ in $L^{2}(\mathbb{R}^{N})$ for all $v\in L^{2}(\mathbb{R}^{N})$. In particular, it holds for $v=\frac{\partial u_{0}}{\partial{x_{i}}}$. We also have $\rho_{m}*u_{0}\to u_{0}$ in $L^{q}(\mathbb{R}^{N})$, as $m\to\infty$. Then we fix $m_{0}\in\mathbb{N}$ sufficiently large so that
$\left\|\rho_{m_{0}}*\frac{\partial u_{0}}{\partial{x_{i}}}-\frac{\partial u_{0}}{\partial{x_{i}}}\right\|_{L^{2}}<\frac{\varepsilon}{10N}$ (6.5)
for all $i\in\{1,\cdots,N\}$, and, for $\psi_{0}:=\rho_{m_{0}}*u_{0}$ we have $supp(\psi_{0})\subset B_{2k_{0}+\sigma}$ and
$\|\psi_{0}-u_{0}\|_{L^{q}}<\frac{\varepsilon}{10n_{0}}.$ (6.6)
Notice that, by the definition of $u_{0}$ and $T_{k}$, we get
$|\psi_{0}(x)|\leq\int_{\mathbb{R}^{N}}\rho_{m_{0}}(x-y)|u_{0}(y)|dy\leq k_{0}\int_{\mathbb{R}^{N}}\rho_{m_{0}}(x-y)dy=k_{0}.$
Thus, using (6.6) we see that
$\|\psi_{0}\chi_{B_{2k_{0}}}-u_{0}\|_{L^{q}_{b}(\{w\leq n_{0}^{q}\})}=\left(\int_{B_{2k_{0}}\cap\{w\leq n_{0}^{q}\}}|\psi_{0}-u_{0}|^{q}w(x)dx\right)^{\frac{1}{q}}\leq n_{0}\|\psi_{0}-u_{0}\|_{L^{q}(B_{2k_{0}})}<\frac{\varepsilon}{10}.$ (6.7)
Also, using (6.4) we have
$\|\psi_{0}-\psi_{0}\chi_{B_{2k_{0}}}\|_{L^{q}_{b}}=\left(\int_{\mathbb{R}^{N}}|\psi_{0}(1-\chi_{B_{2k_{0}}})|^{q}w(x)dx\right)^{\frac{1}{q}}\leq k_{0}\|w\|_{L^{1}(B_{2k_{0}+\sigma}\setminus B_{2k_{0}})}<\frac{\varepsilon}{10}.$ (6.8)
So, due to (6.7), (6.8) and (6.3) we obtain
$\|\psi_{0}-u_{0}\|_{L^{q}_{b}}$ $\leq\|\psi_{0}-\psi_{0}\chi_{B_{2k_{0}}}\|_{L^{q}_{b}}+\|\psi_{0}\chi_{B_{2k_{0}}}-u_{0}\|_{L^{q}_{b}(\{w\leq n_{0}^{q}\})}+\|(\psi_{0}\chi_{B_{2k_{0}}}-u_{0})(1-\chi_{\{w\leq n_{0}^{q}\}})\|_{L^{q}_{b}}$
$<\frac{\varepsilon}{10}+\frac{\varepsilon}{10}+2k_{0}\|(1-\chi_{\{w\leq n_{0}^{q}\}})w\|_{L^{1}(B_{2k_{0}})}<\frac{2\varepsilon}{5}.$ (6.9)
Then, due to (6.5) and (6) it holds
$\|\psi_{0}-u_{0}\|_{E^{q}_{b}}\leq\|\psi_{0}-u_{0}\|_{L^{q}_{b}}+\sum_{i=1}^{N}\left\|\rho_{m_{0}}*\frac{\partial u_{0}}{\partial x_{i}}-\frac{\partial u_{0}}{\partial x_{i}}\right\|_{L^{2}}<\frac{\varepsilon}{2}.$
Finally, using (6.2) we obtain
$\|\psi_{0}-u\|_{E^{q}_{b}}\leq\|\psi_{0}-u_{0}\|_{E^{q}_{b}}+\|u_{0}-u\|_{E^{q}_{b}}<{\varepsilon},$
and conclude this proof. ∎
Acknowledgements
Elisandra Gloss and Bruno Ribeiro would like to thank Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) for the support; in particular, for the grants 201454/2024-6 and 443594/2023-6 (E.G.), grants 314111/2023-9, 443594/2023-6, 201452/2024-3 (B.R.).
Conflict of interest statement
The authors have no relevant financial or non-financial interests to disclose.